### v1

# Partial Condition Numbers for Double Saddle Point Problems

Abstract
This paper presents a unified framework for investigating the partial condition number (CN) of the solution of double saddle point problems (DSPPs) and provides closed-form expressions for it. This unified framework encompasses the well-known partial normwise CN (NCN), partial mixed CN (MCN) and partial componentwise CN (CCN) as special cases. Furthermore, we derive sharp upper bounds for the partial NCN, MCN and CCN, which are computationally efficient and free of expensive Kronecker products. By applying perturbations that preserve the structure of the block matrices of the DSPPs, we analyze the structured partial NCN, MCN and CCN when the block matrices exhibit linear structures. By leveraging the relationship between DSPP and equality constrained indefinite least squares (EILS) problems, we recover the partial CNs for the EILS problem. Numerical results confirm the sharpness of the derived upper bounds and demonstrate their effectiveness in estimating the partial CNs.
keywords:
Partial condition number, Perturbation analysis, Double saddle point problems, Equality constrained indefinite least squares problems, Structured perturbations
pacs:
MSC Classification]15A12, 65F20, 65F35, 65F99
1 Introduction
We consider the following linear system with the double saddle point structure:
$\mathfrak{B}\bm{w}=\bf{b},\leavevmode\nobreak$ (1.1)
where
$\mathfrak{B}=\left[\begin{matrix}A&B^{\bm{\top}}&\bf 0\\ B&-D&C^{\bm{\top}}\\ \bf 0&C&E\end{matrix}\right];\leavevmode\nobreak\ \bm{w}=\left[\begin{matrix}\bm{x}\\ \bm{y}\\ \bm{z}\end{matrix}\right];\leavevmode\nobreak\ \bf{b}=\left[\begin{matrix}\mathbf{b}_{1}\\ \mathbf{b}_{2}\\ \mathbf{b}_{3}\end{matrix}\right];$ (1.2)
$A\in{\mathbb{R}}^{n\times n},$ $B\in{\mathbb{R}}^{m\times n},$ $C\in{\mathbb{R}}^{p\times m},$ $D\in{\mathbb{R}}^{m\times m},$ $E\in{\mathbb{R}}^{p\times p},$ $\mathbf{b}\in{\mathbb{R}}^{\bm{l}\times\bm{l}}$ and $\bm{l}=n+m+p.$ Then the linear system of the form (1.1) is known as the double saddle point problem (DSPP). We refer $\mathfrak{B}$ as the double saddle point matrix. Conditions on the invetability for the matrix $\mathfrak{B}$ have been studied in [1, 2]. To ensure a unique solution to (1.1), throughout the paper, we assume that $\mathfrak{B}$ is nonsingular.
Linear systems of the form (1.1) emerge in numerous computational science and technological applications, such as in finite element discretizations of coupled Stokes-Darcy flow equations [3], the equality constrained indefinite least square problem (EILS) [4], computational fluid dynamics [5], liquid crystal director modeling [6], PDE-constraint optimization problems [7], and so on. Due to the versatile applicability of the DSPP (1.1), their numerical solutions have garnered significant interest and have been extensively investigated in recent years. Several iterative methods and preconditioning techniques have been developed in recent times for solving DSPPs; see [2, 8, 9, 10] and reference therein.
Perturbation analysis is one of the most useful tool in numerical analysis for assessing the sensitivity and stability of the solution of a problem obtained through numerical methods [11]. The condition number (CN) quantifies the worst-case sensitivity of a solution to a small perturbation in the input data matrices. Additionally, CNs are vital in providing first-order error estimates for the relative error of a numerical solution. The general framework of CNs was first introduced by Rice [12]. It primarily focuses on the normwise CN (NCN), where norms are used to quantify both the input perturbations and the resulting errors in the output data. However, a significant limitation of NCN is its inability to take into account the inherent structure of poorly scaled or sparse input data. To overcome this limitation, the mixed CN (MCN) and componentwise CN (CCN) have gained increasing attention in the literature [13, 14, 15]. The former quantifies input perturbations componentwise and output errors using norms, while the later measures both input perturbations and output errors componentwise.
Perturbation analysis and CNs for standard saddle point problems (SPPs) have been extensively studied in the literature; see [16, 17, 18, 19]. Although, these studies fail to leverage the three-by-three block structure of the coefficient matrix $\mathfrak{B}$. Moreover, in many applications, the solution components $\bm{x}$, $\bm{y}$ and $\bm{z}$ of $\bm{w}$ represent different physical entities. For instance, in PDE-constrained optimization problems [7], $\bm{x}$ represents the desired state, $\bm{z}$ corresponds to the control variable, and $\bm{y}$ denotes the Lagrange multipliers. However, the sensitivity of each solution component can vary; hence, analyzing the sensitivity of the individual solution components $\bm{x},$ $\bm{y}$ and $\bm{z}$ or each solution component of $\bm{w}$ becomes crucial [20]. The traditional CNs lack the ability to reveal the conditioning of a specific part of the solution. To tackle this situation, CN of a linear function of the solution has been investigated in the literature, by introducing a matrix $\mathrm{L}$. This matrix enables us to select any particular solution component. The CN of a linear function of the solution is referred to as the partial CN (or the projected CN). For the first time, this class of CN was investigated in [20] for the system of linear equations and later extensively studied for various problems in recent years, for instance, in linear least squares problems [21], weighted least squares problems [22], the indefinite least squares problems [23, 24], total least squares problems [25], generalized SPPs [19].
In recent years, the structured CNs of various problems have been studied, emphasizing the preservation of the linear structure of the original matrices in the perturbation matrices; see, for example [26, 27, 28, 19]. The block matrices $A$, $D$ and $E$ often exhibit particular linear structures in various applications; see [7, 6]. This makes it compelling to explore the structured partial CNs for the DSPP (1.1) by preserving the linear structures of the diagonal block matrices to their corresponding perturbation matrices.
[…]
The key contributions of the paper are highlighted as follows:
  • In this work, we explore a general form of partial CN, refer as unified partial CN, which has a versatile nature and provides a comprehensive framework encompassing the partial NCN, MCN and CCN of the solution of the DSPP.
  • By considering structure-preserving perturbations on $A,$ $D$, and $E,$ when they retain some linear structures, we derive structured partial CNs for the DSPP.
  • By exploring the connection between DSPPs and EILS problems, we demonstrate that our derived CN formula can be used to recover the partial CNs for the EILS problem.
  • Numerical experiments demonstrate that the derived upper bounds provide sharp estimates of the partial CNs. Furthermore, the partial CNs offer precise estimates of the relative forward error in the solution.
The structure of the rest of the paper is as follows. Section 2 introduces a few notations, basic definitions, and preliminaries. Section 3 presents a unified framework partial CN of the solution of the DSPP (1.1). Section 4 focuses on the investigation of the structured partial CNs for the DSPP. In Section 5, we discuss the partial CNs for EILS problems. Section 6 consists of some numerical examples. Section 7 includes the concluding statements.
2 Notation and preliminaries
2.1 Notation
Throughout the paper, ${\mathbb{R}}^{m\times n}$ represents the set of all $m\times n$ real matrices. The symbols $\|\cdot\|_{F}$, $\|\cdot\|_{2},$ and $\|\cdot\|_{\infty}$ stand for the Frobenius norm, spectral norm, and infinity norm of the argument, respectively. The notation ${\bf 1}_{n}\in{\mathbb{R}}^{n}$ represents a vector with all entries equal to $1$. For any vector $z=[z_{1},z_{2},\ldots,z_{n}]^{\bm{\top}}\in{\mathbb{R}}^{n},$ $\Theta_{z}\in{\mathbb{R}}^{n\times n}$ is the diagonal matrix with $i$-th diagonal entry $z_{i}.$ Following [29], for any vector $z\in{\mathbb{R}}^{n},$ we define
$z^{\ddagger}=[z^{\ddagger}_{1},z^{\ddagger}_{2},\ldots,z^{\ddagger}_{n}]^{\bm{\top}},$ (2.1)
where
$z_{i}^{\ddagger}=\left\{\begin{array}[]{cc}\frac{1}{z_{i}},&z_{i}\neq 0,\\ 1,&z_{i}=0.\end{array}\right.$ (2.4)
Moreover, the entrywise division of two vector $z,w\in{\mathbb{R}}^{n}$ is defined as follows:
$\frac{z}{w}=\Theta_{w^{\ddagger}}z.$
For the matrices $X=[x_{ij}]\in{\mathbb{R}}^{m\times n}$ and $Y=[y_{ij}]\in{\mathbb{R}}^{m\times n},$ we define $X\odot Y:=[x_{ij}y_{ij}]\in{\mathbb{R}}^{m\times n}.$ Note that, for $z,w\in{\mathbb{R}}^{n},$ ${\displaystyle{\frac{z}{w}}}=w^{\ddagger}\odot z.$ For $X\in{\mathbb{R}}^{m\times n},$ we set $|X|:=[|x_{ij}|].$ For $X,\,Y\in{\mathbb{R}}^{m\times n},$ the notation $|X|\leavevmode\nobreak\ \leq|Y|$ represents $|x_{ij}|\leavevmode\nobreak\ \leq|y_{ij}|$ for all $1\leq i\leq m$ and $1\leq j\leq n$. For $X=[\bm{x}_{1},\bm{x}_{2},\ldots,\bm{x}_{n}]\in{\mathbb{R}}^{m\times n},$ where $\bm{x}_{i}\in{\mathbb{R}}^{m},$ $i=1,2,\ldots,n,$ we define $\mathrm{v}\mathrm{e}\mathrm{c}(X):=[\bm{x}_{1}^{T},\bm{x}_{2}^{T},\ldots,\bm{x}_{n}^{T}]^{T}\in{\mathbb{R}}^{mn}$. For given matrices $A_{1},A_{2},\ldots,A_{n}$, we use
$\mathrm{v}\mathrm{e}\mathrm{c}(\mathbf{X}):=[\mathrm{v}\mathrm{e}\mathrm{c}(A_{1})^{\bm{\top}},\leavevmode\nobreak\ \mathrm{v}\mathrm{e}\mathrm{c}(A_{2})^{\bm{\top}},\leavevmode\nobreak\ \ldots,\leavevmode\nobreak\ \mathrm{v}\mathrm{e}\mathrm{c}(A_{n})^{\bm{\top}}]^{\bm{\top}},$
where $\mathbf{X}=(A_{1},A_{2},\ldots,A_{n}).$ The Kronecker product [30] of $X\in{\mathbb{R}}^{m\times n}$ and $Y\in{\mathbb{R}}^{p\times q}$ is defined as $X\otimes Y:=[x_{ij}Y]\in{\mathbb{R}}^{mp\times nq}.$ For matrices $A,B,X,Y$ and $Z$ with compatible dimensions, we have the following properties [30, 31]:
$\left\{\begin{array}[]{c}\mathrm{v}\mathrm{e}\mathrm{c}(AZB)=(B^{\bm{\top}}\otimes A)\mathrm{v}\mathrm{e}\mathrm{c}(Z),\\ (X\otimes Y)^{\bm{\top}}=X^{\bm{\top}}\otimes Y^{\bm{\top}},\\ |X\otimes Y|=|X|\otimes|Y|,\\ (A\otimes B)(X\otimes Y)=AX\otimes BY.\\ \end{array}\right.$ (2.9)
2.2 Preliminaries
At the beginning of this subsection, we introduce the concept of the general CN, referred to as the unified CN.
Definition 1.
[32] Let $\Upsilon:{\mathbb{R}}^{p}\mapsto{\mathbb{R}}^{q}$ be a continuous mapping defined on an open set $\Omega_{\Upsilon}\subseteq{\mathbb{R}}^{p}.$ Then, the unified CN of $\Upsilon$ at $\bm{v}\in\Omega_{\Upsilon}$ is defined by
$\mathfrak{K}_{\Upsilon}(\bm{v})=\lim_{\bm{\epsilon}\rightarrow 0}\sup_{0<\|\chi^{\ddagger}\odot{\Delta}\bm{v}\|_{\tau}\leq\bm{\epsilon}}\frac{\left\|\xi^{\ddagger}\odot\left(\Upsilon(\bm{v}+{\Delta}\bm{v})-\Upsilon(\bm{v})\right)\right\|_{\gamma}}{\|\chi^{\ddagger}\odot{\Delta}\bm{v}\|_{\tau}},$ (2.10)
where $\xi\in{\mathbb{R}}^{q},$ ${\chi}\in{\mathbb{R}}^{p}$ are the parameters such that if some entry of $\chi$ is zero, then the corresponding entry of ${\Delta}\bm{v}$ must be zero, and $\|\,\cdot\,\|_{\tau}$ and $\|\,\cdot\,\|_{\gamma}$ are two vector norms defined on ${\mathbb{R}}^{p}$ and ${\mathbb{R}}^{q},$ respectively.
The parameters $\chi$ and $\xi$ in Definition 1 can also be positive real numbers, rather than vectors. Note that, Definition 1 leads to the following bound:
$\left\|\xi^{\ddagger}\odot(\Upsilon(\bm{v}+{\Delta}\bm{v})-\Upsilon(\bm{v}))\right\|_{\gamma}\leq\mathfrak{K}_{\Upsilon}(\bm{v})\|\chi^{\ddagger}\odot{\Delta}\bm{v}\|_{\tau}+\mathcal{O}(\|\chi^{\ddagger}\odot{\Delta}\bm{v}\|^{2}_{\tau}).$ (2.11)
Therefore, the forward error in the solution can be estimated using CNs.
Remark 1.
The unified CN described in Definition 1 represents a broad generalization of various well-known CNs that have been explored in the literature. For example:
  • NCN: Consider $\tau=\gamma=2,$ $\chi=\|\bm{v}\|_{2}{\bf 1}_{p}$ $\in{\mathbb{R}}^{p}$ with $\bm{v}\neq\bf 0$ and $\xi=\|\Upsilon(\bm{v})\|_{2}{\bf 1}_{q}\in{\mathbb{R}}^{q}$ with $\Upsilon(\bm{v})\neq\bf 0,$ then we obtain the NCN, denoted by $\mathfrak{K}^{(2)}_{\Upsilon}(\bm{v}).$
  • MCN: Consider $\tau=\gamma=\infty,$ $\chi=\bm{v}\neq\bf 0,$ $\xi=\|\Upsilon(\bm{v})\|_{\infty}{\bf 1}_{q}\in{\mathbb{R}}^{q}$ with $\Upsilon(\bm{v})\neq\bf 0,$ then we obtain the MCN, denoted by $\mathfrak{K}^{\infty}_{mix,\Upsilon}(\bm{v}).$
  • CCN: Consider $\tau=\gamma=\infty,$ $\chi=\bm{v}\neq\bf 0,$ and $\xi=\Upsilon(\bm{v})\in{\mathbb{R}}^{q}$ with $\Upsilon(\bm{v})\neq\bf 0,$ then we obtain the CCN, denoted by $\mathfrak{K}^{\infty}_{com,\Upsilon}(\bm{v}).$
Next, we present a key result that is essential for the following sections. To derive this, let $\mathbf{H}=(A,B,C,D,E)$ and we set
$\mathrm{v}\mathrm{e}\mathrm{c}(\mathbf{H})=[\mathrm{v}\mathrm{e}\mathrm{c}(A)^{\bm{\top}},\mathrm{v}\mathrm{e}\mathrm{c}(B)^{\bm{\top}},\mathrm{v}\mathrm{e}\mathrm{c}(C)^{\bm{\top}},\mathrm{v}\mathrm{e}\mathrm{c}(D)^{\bm{\top}},\mathrm{v}\mathrm{e}\mathrm{c}(E)^{\bm{\top}}]^{\bm{\top}}.$
Consider ${\Delta}A,{\Delta}B,{\Delta}C,{\Delta}D,{\Delta}E,$ and ${\Delta}\mathbf{b}=[{\Delta}{\bf b}_{1}^{{\bm{\top}}},{\Delta}{\bf b}_{2}^{{\bm{\top}}},{\Delta}{\bf b}_{3}^{{\bm{\top}}}]^{{\bm{\top}}}$ are the perturbations on $A,B,C,D,E,$ and $\mathbf{b},$ respectively. Further, we denote
${\Delta}\mathfrak{B}=\left[\begin{matrix}{\Delta}A&{\Delta}B^{\bm{\top}}&\bf 0\\ {\Delta}B&-{\Delta}D&{\Delta}C^{\bm{\top}}\\ \bf 0&{\Delta}C&{\Delta}E\end{matrix}\right],$
and assume that $\|{\Delta}\mathfrak{B}\|_{2}\leq\bm{\epsilon}\|\mathfrak{B}\|_{2}$ and $\|{\Delta}{\bf b}\|_{2}\leq\bm{\epsilon}\|{\bf b}\|_{2}$. Then, we have the following perturbed DSSP:
$\left[\begin{matrix}A+{\Delta}A&(B+{\Delta}B)^{\bm{\top}}&\bf 0\\ B+{\Delta}B&-(D+{\Delta}D)&(C+{\Delta}C)^{\bm{\top}}\\ \bf 0&C+{\Delta}C&E+{\Delta}E\end{matrix}\right]\left[\begin{matrix}\bm{x}+{\Delta}\bm{x}\\ \bm{y}+{\Delta}\bm{y}\\ \bm{z}+{\Delta}\bm{z}\end{matrix}\right]=\left[\begin{matrix}\mathbf{b}_{1}+{\Delta}\mathbf{b}_{1}\\ \mathbf{b}_{2}+{\Delta}\mathbf{b}_{2}\\ \mathbf{b}_{3}+{\Delta}\mathbf{b}_{3}\end{matrix}\right],$ (2.12)
which has the unique solution $\left[\begin{matrix}\bm{x}+{\Delta}\bm{x}\\ \bm{y}+{\Delta}\bm{y}\\ \bm{z}+{\Delta}\bm{z}\end{matrix}\right]$ when $\|\mathfrak{B}^{-1}\|_{2}\|\Delta\mathfrak{B}\|_{2}<1.$
Consequently, we obtain the following important result.
Lemma 2.1.
Suppose $[\bm{x}^{{\bm{\top}}},\bm{y}^{{\bm{\top}}},\bm{z}^{{\bm{\top}}}]$ and $[(\bm{x}+{\Delta}\bm{x})^{{\bm{\top}}},(\bm{y}+{\Delta}\bm{y})^{{\bm{\top}}},(\bm{z}+{\Delta}\bm{z})^{{\bm{\top}}}]^{{\bm{\top}}}$ are the unique solutions of the original DSPP (1.1) and perturbed DSPP (2.12), respectively. Then, the first-order perturbation expression of $\left[\begin{matrix}{\Delta}\bm{x}^{{\bm{\top}}},{\Delta}\bm{y}^{{\bm{\top}}},{\Delta}\bm{z}^{{\bm{\top}}}\end{matrix}\right]^{{\bm{\top}}}$ is given by
$\left[\begin{matrix}{\Delta}\bm{x}\\ {\Delta}\bm{z}\\ {\Delta}\bm{y}\end{matrix}\right]=-\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}({\Delta}\mathbf{H})\\ {\Delta}{\bf b}\end{matrix}\right]+\mathcal{O}(\bm{\epsilon}^{2}),$ (2.13)
where
$\mathcal{G}=\left[\begin{matrix}\bm{x}^{\bm{\top}}\otimes I_{n}&I_{n}\otimes\bm{y}^{\bm{\top}}&\bf 0&\bf 0&\bf 0\\ \bf 0&\bm{x}^{\bm{\top}}\otimes I_{m}&I_{m}\otimes\bm{z}^{\bm{\top}}&-\bm{y}^{\bm{\top}}\otimes I_{m}&\bf 0\\ \bf 0&\bf 0&\bm{y}^{\bm{\top}}\otimes I_{p}&\bf 0&\bm{z}^{\bm{\top}}\otimes I_{p}\end{matrix}\right]\in{\mathbb{R}}^{\bm{l}\times\bm{s}},$ (2.14)
$\mathrm{v}\mathrm{e}\mathrm{c}({\Delta}\mathbf{H})=[\mathrm{v}\mathrm{e}\mathrm{c}({\Delta}A)^{\bm{\top}},\mathrm{v}\mathrm{e}\mathrm{c}({\Delta}B)^{\bm{\top}},\mathrm{v}\mathrm{e}\mathrm{c}({\Delta}C)^{\bm{\top}},\mathrm{v}\mathrm{e}\mathrm{c}({\Delta}D)^{\bm{\top}},\mathrm{v}\mathrm{e}\mathrm{c}({\Delta}E)^{\bm{\top}}]^{\bm{\top}},$
and ${\bm{s}}=(n^{2}+m^{2}+p^{2}+nm+mp).$
Proof.
Combining (1.1) and (2.12), we obtain
$\left[\begin{matrix}A&B^{\bm{\top}}&\bf 0\\ B&-D&C^{\bm{\top}}\\ \bf 0&C&E\end{matrix}\right]\left[\begin{matrix}{\Delta}\bm{x}\\ {\Delta}\bm{y}\\ {\Delta}\bm{z}\end{matrix}\right]=\left[\begin{matrix}{\Delta}\mathbf{b}_{1}\\ {\Delta}\mathbf{b}_{2}\\ {\Delta}\mathbf{b}_{3}\end{matrix}\right]-\left[\begin{matrix}{\Delta}A\bm{x}+{\Delta}B^{\bm{\top}}\bm{y}\\ {\Delta}B\bm{x}-{\Delta}D\bm{y}+{\Delta}C^{\bm{\top}}\bm{z}\\ {\Delta}C\bm{y}+{\Delta}E\bm{z}\end{matrix}\right]+\mathcal{O}(\bm{\epsilon}^{2}).$ (2.15)
Thus, the proof follows by applying the vec operator on (2.15) and utilizing the properties listed in (2.9). $\blacksquare$
3 Partial Unified CNs for the DSPP
This section primarily focuses on developing a unified framework for the partial CN for the solution $\bm{w}=[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}$ of the DSPP (1.1). As special cases, we also derive the compact formulae and computationally efficient upper bounds for the partial NCN, MCN and CCN.
To derive the partial unified CN of the DSPP (1.1), we define the following mapping:
$\bm{\varphi}:\,$ ${\mathbb{R}}^{n\times n}\times{\mathbb{R}}^{m\times n}\times{\mathbb{R}}^{p\times m}\times{\mathbb{R}}^{m\times m}\times{\mathbb{R}}^{p\times p}\times{\mathbb{R}}^{\bm{l}}\rightarrow{\mathbb{R}}^{k}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \bm{\varphi}(\mathbf{H},\mathbf{b})=\mathrm{L}\left[\begin{matrix}\bm{x}\\ \bm{y}\\ \bm{z}\end{matrix}\right]=\mathrm{L}\mathfrak{B}^{-1}\mathbf{b},$ (3.1)
where $\mathrm{L}\in{\mathbb{R}}^{k\times\bm{l}}(k\leq\bm{l}).$ Following the Definition 1, we now define the partial unified CN for the DSPP using the mapping $\bm{\varphi}$ as follows.
Definition 2.
Suppose $\bm{w}=[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}$ is the unique solution of the DSPP (1.1) and $\mathrm{L}\in{\mathbb{R}}^{k\times\bm{l}}.$ Consider the map $\bm{\varphi}$ defined as in (3). Then, the partial unified CN of $\bm{w}=[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}$ with respect to (w.r.t.) $\mathrm{L}$ is defined as follows:
$\mathfrak{K}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}):=\lim_{\bm{\epsilon}\rightarrow 0}\sup_{0<\left\|\mathrm{v}\mathrm{e}\mathrm{c}\left(\Psi^{\ddagger}\odot\Delta\mathbf{H},\,\bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\right)\right\|_{\tau}\leq\bm{\epsilon}}\frac{\left\|\xi^{\ddagger}_{\mathrm{L}}\odot\left(\bm{\varphi}(\mathbf{H}+{\Delta}\mathbf{H},\mathbf{b}+{\Delta}\mathbf{b})-\bm{\varphi}(\mathbf{H},{\bf b})\right)\right\|_{\gamma}}{\left\|\mathrm{v}\mathrm{e}\mathrm{c}\left(\Psi^{\ddagger}\odot\Delta\mathbf{H},\bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\right)\right\|_{\tau}},$
where $\xi_{\mathrm{L}}\in{\mathbb{R}}^{k},$ $\Psi=(\Psi_{A},\Psi_{B},\Psi_{C},\Psi_{D},\Psi_{E}),$ $\Psi_{A}\in{\mathbb{R}}^{n\times n},$ $\Psi_{B}\in{\mathbb{R}}^{m\times n},$ $\Psi_{C}\in{\mathbb{R}}^{p\times m},$ $\Psi_{D}\in{\mathbb{R}}^{m\times m},$ $\Psi_{E}\in{\mathbb{R}}^{p\times p}$ and $\bm{\chi}\in{\mathbb{R}}^{\bm{l}}$ are the parameters with the assumptions that if some entries of $\Psi$ and $\bm{\chi}$ are zero, then the corresponding entry of ${\Delta}\mathbf{H}$ and ${\Delta}{\bf b},$ respectively, must be zero.
Remark 2.
In the context of Remark 1, to obtain the partial NCN, we consider $\xi_{\mathrm{L}}=\|\mathrm{L}[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{{\bm{\top}}}\|_{2}{\bf 1}_{\bm{l}},$ for the partial MCN, we consider $\xi_{\mathrm{L}}=\|\mathrm{L}[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{{\bm{\top}}}\|_{\infty}{\bf 1}_{\bm{l}},$ and for partial CCN, we consider $\xi_{\mathrm{L}}=\mathrm{L}[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{{\bm{\top}}}.$
In the following theorem, we provide compact and closed-form expression for the partial unified CN.
Theorem 1.
Suppose $\bm{w}=[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}$ is the unique solution of the DSPP (1.1) and $\mathrm{L}\in{\mathbb{R}}^{k\times\bm{l}}.$ Then, the partial unified CN of $\bm{w}=[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}$ w.r.t. $\mathrm{L}$ is given by
$\mathfrak{K}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})=\left\|\Theta_{\xi^{\ddagger}_{\mathrm{L}}}\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\Theta_{\mathrm{v}\mathrm{e}\mathrm{c}(\Psi)}&\bf 0\\ \bf 0&\Theta_{\bm{\chi}}\end{matrix}\right]\right\|_{\tau,\gamma},$ (3.2)
where $\|\cdot\leavevmode\nobreak\ \|_{\tau,\gamma}$ is the matrix norm induced by vector norms $\|\cdot\leavevmode\nobreak\ \|_{\tau}$ and $\|\cdot\leavevmode\nobreak\ \|_{\gamma}.$
Proof.
From the definition of the mapping $\bm{\varphi}$ in (3) and Lemma 2.1, we get
$\bm{\varphi}(\mathbf{H}+{\Delta}\mathbf{H},{\bf b}+{\Delta}{\bf b})-\bm{\varphi}(\mathbf{H},{\bf b})$ $=\mathrm{L}\left[\begin{matrix}\bm{x}+{\Delta}\bm{x}\\ \bm{y}+{\Delta}\bm{y}\\ \bm{z}+{\Delta}\bm{z}\end{matrix}\right]-\mathrm{L}\left[\begin{matrix}\bm{x}\\ \bm{y}\\ \bm{z}\end{matrix}\right]$
$=\mathrm{L}\left[\begin{matrix}{\Delta}\bm{x}\\ {\Delta}\bm{y}\\ {\Delta}\bm{z}\end{matrix}\right]$
$=-\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}({\Delta}\mathbf{H})\\ {\Delta}{\bf b}\end{matrix}\right]+\mathcal{O}(\bm{\epsilon}^{2}).$ (3.3)
By considering the requirement on $\Psi$ and $\bm{\chi},$ we have
$\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}({\Delta}\mathbf{H})\\ {\Delta}{\bf b}\end{matrix}\right]=\left[\begin{matrix}\Theta_{\mathrm{v}\mathrm{e}\mathrm{c}(\Psi)}&\bf 0\\ \bf 0&\Theta_{\bm{\chi}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right].$ (3.4)
Substituting (3.4) into (3.3) and from Definition 2, we obtain
$\mathfrak{K}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $=\sup_{\left\|\mathrm{v}\mathrm{e}\mathrm{c}\left(\Psi^{\ddagger}\odot\Delta\mathbf{H},\,\bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\right)\right\|_{\tau}\neq 0}\frac{\left\|\Theta_{\xi^{\ddagger}_{\mathrm{L}}}\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\Theta_{\mathrm{v}\mathrm{e}\mathrm{c}(\Psi)}&\bf 0\\ \bf 0&\Theta_{\bm{\chi}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]\right\|_{\gamma}}{\left\|\mathrm{v}\mathrm{e}\mathrm{c}\left(\Psi^{\ddagger}\odot\Delta\mathbf{H},\bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\right)\right\|_{\tau}}$
$=\left\|\Theta_{\xi^{\ddagger}_{\mathrm{L}}}\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\Theta_{\mathrm{v}\mathrm{e}\mathrm{c}(\Psi)}&\bf 0\\ \bf 0&\Theta_{\bm{\chi}}\end{matrix}\right]\right\|_{\tau,\gamma}.$ (3.5)
Hence, the proof is completed. $\blacksquare$
Next, we derive various partial CNs by considering specific norms. In the following result, we focus on when $\tau=\gamma=2$.
Theorem 2.
Consider $\tau=\gamma=2$ and assuming that $\Psi,$ $\bm{\chi}$ and $\xi_{\mathrm{L}}$ are positive real numbers, then the partial CN has the following forms:
$\mathfrak{K}^{(2)}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})=\frac{\left\|\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\Psi\mathcal{G}&-\bm{\chi}I_{\bm{l}}\end{matrix}\right]\right\|_{2}}{\xi_{\mathrm{L}}},$ (3.6)
$\widehat{\mathfrak{K}}^{(2)}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})=\frac{\left\|\mathrm{L}\mathfrak{B}^{-1}({\Psi}^{2}\mathcal{J}+{\bm{\chi}}^{2}I_{\bm{l}})(\mathfrak{B}^{-1})^{{\bm{\top}}}\mathrm{L}^{\bm{\top}}\right\|^{1/2}_{2}}{\xi_{\mathrm{L}}},$ (3.7)
where $\mathcal{J}\in{\mathbb{R}}^{\bm{l}\times\bm{l}}$ is given by
$\mathcal{J}=\left[\begin{matrix}(\|\bm{x}\|_{2}^{2}+\|\bm{y}\|_{2}^{2})I_{n}&\bm{x}\bm{y}^{\bm{\top}}&{\bf 0}\\ \bm{y}\bm{x}^{\bm{\top}}&(\|\bm{x}\|_{2}^{2}+\|\bm{y}\|_{2}^{2}+\|\bm{z}\|_{2}^{2})I_{m}&\bm{y}\bm{z}^{\bm{\top}}\\ \bf 0&\bm{z}\bm{y}^{\bm{\top}}&(\|\bm{y}\|_{2}^{2}+\|\bm{z}\|_{2}^{2})I_{p}\end{matrix}\right].$
Proof.
Since $\tau=\gamma=2$ and $\Psi,$ $\bm{\chi}$ and $\xi_{\mathrm{L}}$ are positive real numbers, from Theorem 1, we obtain
$\mathfrak{K}^{(2)}_{1,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})=\frac{\left\|\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\Psi\mathcal{G}&-\bm{\chi}I_{\bm{l}}\end{matrix}\right]\right\|_{2}}{\xi_{\mathrm{L}}}.$ (3.8)
Using the property that, for $Z\in{\mathbb{R}}^{m\times n},$ $\|Z\|_{2}=\|ZZ^{\bm{\top}}\|_{2}^{1/2},$ we have
$\left\|\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\Psi\mathcal{G}&-\bm{\chi}I_{\bm{l}}\end{matrix}\right]\right\|_{2}$ $=\left\|\mathrm{L}\mathfrak{B}^{-1}(\Psi^{2}\mathcal{G}\mathcal{G}^{\bm{\top}}+\bm{\chi}^{2}I_{\bm{l}})(\mathfrak{B}^{-1})^{{\bm{\top}}}\mathrm{L}^{\bm{\top}}\right\|_{2}^{1/2}$
$=\left\|\mathrm{L}\mathfrak{B}^{-1}(\Psi^{2}\mathcal{J}+\bm{\chi}^{2}I_{\bm{l}})(\mathfrak{B}^{-1})^{{\bm{\top}}}\mathrm{L}^{\bm{\top}}\right\|_{2}^{1/2}.$ (3.9)
Hence, the proof follows by substituting (3) into (3.8). $\blacksquare$
Remark 3.
Notably, the equivalent expression of $\widehat{\mathfrak{K}}^{(2)}_{\Phi}(\mathbf{H},{\bf b};\mathrm{L})$ in (3.7) is free of computationally expensive Kronecker products. Moreover, the matrices in (3.6) and (3.7) have dimensions $k\times(\bm{l+s})$ and $k\times k$, respectively. Hence, the expression in (3.7) significantly reduces the storage requirements.
Remark 4.
The partial CN in Theorem 2 is a simplified version of the partial NCN for the solution of the DSPP (1.1). The NCN for ${\bm{w}}=[\bm{x}^{{\bm{\top}}},\bm{y}^{{\bm{\top}}},\bm{z}^{{\bm{\top}}}]^{{\bm{\top}}},\bm{x},\bm{y}$ and $\bm{z}$ can be obtained by considering
$\mathrm{L}=I_{\bm{l}},\leavevmode\nobreak\ \left[\begin{matrix}I_{n}&{\bf 0}_{n\times(m+p)}\end{matrix}\right],\leavevmode\nobreak\ \left[\begin{matrix}{\bf 0}_{m\times n}&I_{m}&{\bf 0}_{m\times p}\end{matrix}\right]\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \left[\begin{matrix}{\bf 0}_{p\times(n+m)}&I_{p}\end{matrix}\right],$
respectively, in Theorem 2.
In the next result, we provide an easily computable upper bound for the partial CN $\mathfrak{K}^{(2)}_{\Phi}(\mathbf{H},{\bf b};\mathrm{L}).$
Corollary 3.1.
Under the assumption of Theorem 2, we have the following upper bound:
$\mathfrak{K}^{(2)}_{1,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})\leq\mathfrak{K}^{(2),u}_{1,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}):=\frac{\|\mathrm{L}\mathfrak{B}^{-1}\|_{2}}{\xi_{\mathrm{L}}}\left(\Psi\|\mathcal{J}\|^{1/2}_{2}+\bm{\chi}\right).$ (3.10)
Proof.
Using the properties of the spectral norm that for the matrices $X$ and $Y$ of appropriate sizes, $\left\|\left[\begin{matrix}X&Y\end{matrix}\right]\right\|_{2}\leq\|X\|_{2}+\|Y\|_{2}$ and $\|XY\|_{2}\leq\|X\|_{2}\|Y\|_{2},$ and from (3.6), we obtain
$\left\|\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\Psi\mathcal{G}&-\bm{\chi}I_{\bm{l}}\end{matrix}\right]\right\|_{2}$ $\leq\Psi\|\mathrm{L}\mathfrak{B}^{-1}\mathcal{G}\|_{2}+\bm{\chi}\|\mathrm{L}\mathfrak{B}^{-1}\|_{2}$
$\leq\Psi\|\mathrm{L}\mathfrak{B}^{-1}\|_{2}\|\mathcal{G}\|_{2}+\bm{\chi}\|\mathrm{L}\mathfrak{B}^{-1}\|_{2}$
$=\Psi\|\mathrm{L}\mathfrak{B}^{-1}\|_{2}\|\mathcal{J}\|^{1/2}_{2}+\bm{\chi}\|\mathrm{L}\mathfrak{B}^{-1}\|_{2}.$ (3.11)
Hence, the proof follows from (3.6) and (3). $\blacksquare$
In the following theorem, we investigate the partial CN for the DSPP when $\tau=\gamma=\infty$, from which we derive the partial MCN and CCN.
Theorem 3.
When $\tau=\gamma=\infty,$ the partial CN is given as follows:
$\mathfrak{K}^{\infty}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})=\left\||\Theta_{\xi^{\ddagger}_{\mathrm{L}}}|\left|\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\right|\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}(|\Psi|)\\ |\bm{\chi}|\end{matrix}\right]\right\|_{\infty}.$ (3.12)
Moreover, we consider $\Psi=\mathbf{H},$ $\bm{\chi}={\bf b}.$ Then, if we set $\xi_{\mathrm{L}}=\|\mathrm{L}[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}\|_{\infty}{\bf{1}}_{\bm{l}},$ the partial MCN is given by
$\mathfrak{K}^{\infty}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})=\frac{\left\|\left|\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\right|\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}(|\mathbf{H}|)\\ |{\bf b}|\end{matrix}\right]\right\|_{\infty}}{\|\mathrm{L}[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}\|_{\infty}}$ (3.13)
and if we set $\xi_{\mathrm{L}}=\mathrm{L}[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}},$ the partial CCN is given by
$\mathfrak{K}^{\infty}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})=\left\|\frac{\left|\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\right|\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}(|\mathbf{H}|)\\ |{\bf b}|\end{matrix}\right]}{\left|\mathrm{L}[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}\right|}\right\|_{\infty}.$ (3.14)
Proof.
Consider $\tau=\gamma=\infty,$ then from Theorem 1, we have
$\mathfrak{K}^{\infty}_{\Phi}(\mathbf{H},{\bf b};\mathrm{L})$ $=\left\|\Theta_{\xi^{\ddagger}_{\mathrm{L}}}\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\Theta_{\mathrm{v}\mathrm{e}\mathrm{c}(\Psi)}&\bf 0\\ \bf 0&\Theta_{\bm{\chi}}\end{matrix}\right]\right\|_{\infty}$
$=\left\||\Theta_{\xi^{\ddagger}_{\mathrm{L}}}|\left|\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\right|\left[\begin{matrix}|\Theta_{\mathrm{v}\mathrm{e}\mathrm{c}(\Psi)}|&\bf 0\\ \bf 0&|\Theta_{\bm{\chi}}|\end{matrix}\right]\right\|_{\infty}$
$=\left\||\Theta_{\xi^{\ddagger}_{\mathrm{L}}}|\left|\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\right|\left[\begin{matrix}|\Theta_{\mathrm{v}\mathrm{e}\mathrm{c}(\Psi)}|&\bf 0\\ \bf 0&|\Theta_{\bm{\chi}}|\end{matrix}\right]{\bf 1}_{s+\bm{l}}\right\|_{\infty}$
$=\left\||\Theta_{\xi^{\ddagger}_{\mathrm{L}}}|\left|\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\right|\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}(|\Psi|)\\ |\bm{\chi}|\end{matrix}\right]\right\|_{\infty}.$
Rest of the proof followings considering $\Psi=\mathbf{H},$ $\bm{\chi}={\bf b},$ $\xi_{\mathrm{L}}=\|\mathrm{L}[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}\|_{\infty}{\bf{1}}_{\bm{l}}$ or $\xi_{\mathrm{L}}=\mathrm{L}[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}.$blacksquare
Remark 5.
The MCN and CCN for ${\bm{w}}=[\bm{x}^{{\bm{\top}}},\bm{y}^{{\bm{\top}}},\bm{z}^{{\bm{\top}}}]^{{\bm{\top}}},\bm{x},\bm{y}$ and $\bm{z}$ can be obtained by considering
$\mathrm{L}=I_{\bm{l}},\leavevmode\nobreak\ \left[\begin{matrix}I_{n}&{\bf 0}_{n\times(m+p)}\end{matrix}\right],\leavevmode\nobreak\ \left[\begin{matrix}{\bf 0}_{m\times n}&I_{m}&{\bf 0}_{m\times p}\end{matrix}\right]\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \left[\begin{matrix}{\bf 0}_{p\times(n+m)}&I_{p}\end{matrix}\right],$
respectively, in (3.13) and (3.14) of Theorem 3.
In the following result, we provide sharp upper bounds for the partial MCN and CCN obtained in Theorem 3.
Corollary 3.2.
Assume that the conditions in Theorem 3 hold. Then
$\mathfrak{K}^{\infty}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})\leq\mathfrak{K}^{\infty,u}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}):=\frac{\left\||\mathrm{L}\mathfrak{B}^{-1}|\left(|\mathcal{H}|+|{\bf b}|\right)\right\|_{\infty}}{\left\|\mathrm{L}[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}\right\|_{\infty}}$
and
$\mathfrak{K}^{\infty}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})\leq\mathfrak{K}^{\infty,u}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}):=\left\|\frac{|\mathrm{L}\mathfrak{B}^{-1}|\left(|\mathcal{H}|+|{\bf b}|\right)}{\mathrm{L}[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}}\right\|_{\infty},$
where $\mathcal{H}=\left[\begin{matrix}|A||\bm{x}|+|B^{\bm{\top}}||\bm{y}|\\ |B||\bm{x}|+|D^{\bm{\top}}||\bm{y}|+|C^{\bm{\top}}||\bm{z}|\\ |C||\bm{y}|+|E||\bm{z}|\end{matrix}\right].$
Proof.
Utilizing the properties of Kronecker product in (2.9), we have
$\left|\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}\leavevmode\nobreak\ &-I_{\bm{l}}\end{matrix}\right]\right|\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}(|\mathbf{H}|)\\ |{\bf b}|\end{matrix}\right]\leq|\mathrm{L}\mathfrak{B}^{-1}|\left[\begin{matrix}|\mathcal{G}|&I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}(|\mathbf{H}|)\\ |{\bf b}|\end{matrix}\right]$
$=|\mathrm{L}\mathfrak{B}^{-1}|(|\mathcal{G}|\mathrm{v}\mathrm{e}\mathrm{c}(|\mathbf{H}|)+|{\bf b}|)$
$=|\mathrm{L}\mathfrak{B}^{-1}|\left(\left[\begin{matrix}(|\bm{x}|^{\bm{\top}}\otimes I_{n})\mathrm{v}\mathrm{e}\mathrm{c}(|A|)+(I_{n}\otimes|\bm{y}|^{\bm{\top}})\mathrm{v}\mathrm{e}\mathrm{c}(|B|)\\ (|\bm{x}|^{\bm{\top}}\otimes I_{m})\mathrm{v}\mathrm{e}\mathrm{c}(|B|)+(I_{m}\otimes|\bm{z}|^{\bm{\top}})\mathrm{v}\mathrm{e}\mathrm{c}(|C|)+(|\bm{y}|^{\bm{\top}}\otimes I_{m})\mathrm{v}\mathrm{e}\mathrm{c}(|D|)\\ (|\bm{y}|^{\bm{\top}}\otimes I_{p})\mathrm{v}\mathrm{e}\mathrm{c}(|C|)+(|\bm{z}|^{\bm{\top}}\otimes I_{p})\mathrm{v}\mathrm{e}\mathrm{c}(|E|)\end{matrix}\right]+|{\bf b}|\right)$
$=|\mathrm{L}\mathfrak{B}^{-1}|\left(\left|\left[\begin{matrix}|A||\bm{x}|+|B^{\bm{\top}}||\bm{y}|\\ |B||\bm{x}|+|D^{\bm{\top}}||\bm{y}|+|C^{\bm{\top}}||\bm{z}|\\ |C||\bm{y}|+|E||\bm{z}|\end{matrix}\right]\right|+|{\bf b}|\right).$ (3.15)
From (3) and the expressions of partial MCN and CCN in Theorem 3, we get
$\mathfrak{K}^{\infty}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})\leq\mathfrak{K}^{\infty,u}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \mathfrak{K}^{\infty}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})\leq\mathfrak{K}^{\infty,u}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}).$
Hence, the proof follows. $\blacksquare$
4 Structured partial CNs
Consider three subspaces $\mathbb{S}_{1}\subseteq{\mathbb{R}}^{n\times n}$, $\mathbb{S}_{2}\subseteq{\mathbb{R}}^{m\times m}$ and $\mathbb{S}_{3}\subseteq{\mathbb{R}}^{p\times p}$ consisting of three distinct linear structured matrices, such as symmetric. Suppose that the corresponding dimensions of the linear subspaces are $s,$ $r$ and $q,$ respectively. Let $A\in\mathbb{S}_{1},$ $D\in\mathbb{S}_{2}$ and $E\in\mathbb{S}_{3},$ then according to [33, 34, 26], there exist unique generating vectors $\bm{a}\in{\mathbb{R}}^{s}$ $\bm{d}\in{\mathbb{R}}^{r}$ and $\bm{e}\in{\mathbb{R}}^{q}$ such that
$\mathrm{v}\mathrm{e}\mathrm{c}(A)=\bm{\Phi}_{\mathbb{S}_{1}}\bm{a},\leavevmode\nobreak\ \mathrm{v}\mathrm{e}\mathrm{c}(D)=\bm{\Phi}_{\mathbb{S}_{2}}\bm{d}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{v}\mathrm{e}\mathrm{c}(E)=\bm{\Phi}_{\mathbb{S}_{3}}\bm{e},$ (4.1)
where $\bm{\Phi}_{\mathbb{S}_{1}}\in{\mathbb{R}}^{n^{2}\times s},$ $\bm{\Phi}_{\mathbb{S}_{2}}\in{\mathbb{R}}^{m^{2}\times r}$ and $\bm{\Phi}_{\mathbb{S}_{3}}\in{\mathbb{R}}^{p^{2}\times q}.$ These matrices are fixed for each specific structure and encapsulate the information corresponding to the linear structure of their respective subspaces.
Let $\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\mathbf{H})=[\bm{a}^{\bm{\top}},\mathrm{v}\mathrm{e}\mathrm{c}(B)^{\bm{\top}},\mathrm{v}\mathrm{e}\mathrm{c}(C)^{\bm{\top}},\bm{d}^{\bm{\top}},\bm{e}^{\bm{\top}}]^{\bm{\top}}.$ Then the structured partial CN for solution $\bm{w}=[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}$ of the DSPP (1.1) w.r.t. $\mathrm{L}$ is given by
$\mathfrak{K}^{\mathbb{S}}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}):=\lim_{\bm{\epsilon}\rightarrow 0}\sup_{\begin{subarray}{c}0<\left\|\mathrm{v}\mathrm{e}\mathrm{c}\left(\Psi^{\ddagger}\odot\Delta\mathbf{H},\,\bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\right)\right\|_{\tau}\leq\bm{\epsilon}\\ {\Delta}A\in\,\mathbb{S}_{1},\,{\Delta}D\in\,\mathbb{S}_{2},\,{\Delta}E\in\,\mathbb{S}_{3}\end{subarray}}\frac{\left\|\xi^{\ddagger}_{\mathrm{L}}\odot\left(\bm{\varphi}(\mathbf{H}+{\Delta}\mathbf{H},\mathbf{b}+{\Delta}\mathbf{b})-\bm{\varphi}(\mathbf{H},{\bf b})\right)\right\|_{\gamma}}{\left\|\mathrm{v}\mathrm{e}\mathrm{c}\left(\Psi^{\ddagger}\odot\Delta\mathbf{H},\bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\right)\right\|_{\tau}},$ (4.2)
where $\xi_{\mathrm{L}}\in{\mathbb{R}}^{k},$ $\Psi=(\Psi_{A},\Psi_{B},\Psi_{C},\Psi_{D},\Psi_{E}),$ $\Psi_{A}\in\mathbb{S}_{1},$ $\Psi_{B}\in{\mathbb{R}}^{m\times n},$ $\Psi_{C}\in{\mathbb{R}}^{p\times m},$ $\Psi_{D}\in\mathbb{S}_{2},$ $\Psi_{E}\in\mathbb{S}_{3}$ and $\bm{\chi}\in{\mathbb{R}}^{\bm{l}}$.
Since the matrices ${\Delta}A,\Psi_{A}\in\mathbb{S}_{1},$ ${\Delta}D,\Psi_{D}\in\mathbb{S}_{2}$ and ${\Delta}E,\Psi_{E}\in\mathbb{S}_{3},$ as in (4.1), we have
$\mathrm{v}\mathrm{e}\mathrm{c}({\Delta}A)=\bm{\Phi}_{\mathbb{S}_{1}}{\Delta}\bm{a},\leavevmode\nobreak\ \mathrm{v}\mathrm{e}\mathrm{c}(\Psi_{A})=\bm{\Phi}_{\mathbb{S}_{1}}\psi_{A},\leavevmode\nobreak\ \mathrm{v}\mathrm{e}\mathrm{c}({\Delta}D)=\bm{\Phi}_{\mathbb{S}_{2}}{\Delta}\bm{d},$ (4.3)
$\mathrm{v}\mathrm{e}\mathrm{c}(\Psi_{D})=\bm{\Phi}_{\mathbb{S}_{2}}\psi_{D},\leavevmode\nobreak\ \mathrm{v}\mathrm{e}\mathrm{c}({\Delta}E)=\bm{\Phi}_{\mathbb{S}_{3}}{\Delta}\bm{e},\leavevmode\nobreak\ \mathrm{v}\mathrm{e}\mathrm{c}(\Psi_{E})=\bm{\Phi}_{\mathbb{S}_{3}}\psi_{E},$ (4.4)
where ${\Delta}\bm{a},$ ${\Delta}\bm{d},$ ${\Delta}\bm{e},$ $\psi_{D},$ $\psi_{A},$ and $\psi_{E}$ are the unique generating vectors of ${\Delta}A,$ ${\Delta}D,$ ${\Delta}E,$ $\Psi_{A},$ $\Psi_{D},$ and $\Psi_{E}$, respectively. Note that, $\Psi_{A}^{\ddagger}\in\mathbb{S}_{1},$ $\Psi_{D}^{\ddagger}\in\mathbb{S}_{2}$ and $\Psi_{E}^{\ddagger}\in\mathbb{S}_{3}$. Consequently, we obtain
$\mathrm{v}\mathrm{e}\mathrm{c}(\Psi_{A}^{\ddagger}\odot{\Delta}A)$ $=\bm{\Phi}_{\mathbb{S}_{1}}(\psi_{A}^{\ddagger}\odot{\Delta}\bm{a}),\leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{v}\mathrm{e}\mathrm{c}(\Psi_{D}^{\ddagger}\odot{\Delta}D)=\bm{\Phi}_{\mathbb{S}_{2}}(\psi_{D}^{\ddagger}\odot{\Delta}\bm{d})$ (4.5)
$\mathrm{v}\mathrm{e}\mathrm{c}(\Psi_{E}^{\ddagger}\odot{\Delta}E)=\bm{\Phi}_{\mathbb{S}_{3}}(\psi_{E}^{\ddagger}\odot{\Delta}\bm{e}).$ (4.6)
Subsequently, we obtain the following result.
Lemma 4.1.
Let ${\Delta}A,\Psi_{A}\in\mathbb{S}_{1},$ ${\Delta}D,\Psi_{D}\in\mathbb{S}_{2},$ ${\Delta}E,\Psi_{E}\in\mathbb{S}_{3},$ $B,\Psi_{B}\in{\mathbb{R}}^{m\times n},$ $C,\Psi_{C}\in{\mathbb{R}}^{p\times m},$ and ${\bf b},\bm{\chi}\in{\mathbb{R}}^{\bm{l}}.$ Then, we have
$\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}(\Psi^{\ddagger}\odot\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]=\left[\begin{matrix}\bm{\Phi}_{\mathbb{S}}&\bf 0\\ \bf 0&I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot H)\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right],$ (4.7)
where $\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot\mathbf{H})=[(\psi_{A}^{\ddagger}\odot{\Delta}\bm{a})^{\bm{\top}},(\Psi_{B}^{\ddagger}\odot\mathrm{v}\mathrm{e}\mathrm{c}({\Delta}B))^{\bm{\top}},(\Psi_{C}^{\ddagger}\odot\mathrm{v}\mathrm{e}\mathrm{c}({\Delta}C))^{\bm{\top}},(\psi_{D}^{\ddagger}\odot{\Delta}\bm{d})^{\bm{\top}},(\psi_{D}^{\ddagger}\odot{\Delta}\bm{e})^{\bm{\top}}]^{\bm{\top}}$ and
$\bm{\Phi}_{\mathbb{S}}=\left[\begin{matrix}\bm{\Phi}_{\mathbb{S}_{1}}&\bf 0&\bf 0&\bf 0\\ \bf 0&I_{mn+mp}&\bf 0&\bf 0\\ \bf 0&\bf 0&\bm{\Phi}_{\mathbb{S}_{2}}&\bf 0\\ \bf 0&\bf 0&\bf 0&\bm{\Phi}_{\mathbb{S}_{3}}\end{matrix}\right].$
Proof.
The proof follows immediately using identities in (4.5) and (4.6). $\blacksquare$
In the following theorem, we present closed-form expressions for the structured partial CN by considering $\tau=\gamma=2.$
Theorem 4.
Let $A\in\mathbb{S}_{1},$ $D\in\mathbb{S}_{2},$ $E\in\mathbb{S}_{3}$ and $\mathrm{L}\in{\mathbb{R}}^{k\times\bm{l}}.$ Suppose that ${\bm{w}}=[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}$ is the unique solution of DSPP (1.1) and consider $\tau=\gamma=2.$ Then the structured partial CN of ${\bm{w}}=[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}$ w.r.t. $\mathrm{L}$ is given by
$\mathfrak{K}^{(2),\mathbb{S}}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})=\left\|\Theta_{\xi^{\ddagger}_{\mathrm{L}}}\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\Theta_{\mathrm{v}\mathrm{e}\mathrm{c}(\Psi)}\bm{\Phi}_{\mathbb{S}}\mathfrak{D}^{-1}_{\mathbb{S}}&\bf 0\\ \bf 0&\Theta_{\bm{\chi}}\end{matrix}\right]\right\|_{2},$
where
$\mathfrak{D}_{\mathbb{S}}=\left[\begin{matrix}\Theta_{\bm{u}_{1}}&\bf 0&\bf 0&\bf 0\\ \bf 0&I_{mn+mp}&\bf 0&\bf 0\\ \bf 0&\bf 0&\Theta_{\bm{u}_{2}}&\bf 0\\ \bf 0&\bf 0&\bf 0&\Theta_{\bm{u}_{3}}\end{matrix}\right],$
$\bm{u}_{1}=[\|\bm{\Phi}_{\mathbb{S}_{1}}(:,1)\|_{2},\ldots,\|\bm{\Phi}_{\mathbb{S}_{1}}(:,s)\|_{2}]^{\bm{\top}},\leavevmode\nobreak\ \leavevmode\nobreak\ \bm{u}_{2}=[\|\bm{\Phi}_{\mathbb{S}_{2}}(:,1)\|_{2},\ldots,\|\bm{\Phi}_{\mathbb{S}_{2}}(:,r)\|_{2}]^{\bm{\top}},$
$\text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \bm{u}_{3}=[\|\bm{\Phi}_{\mathbb{S}_{3}}(:,1)\|_{2},\ldots,\|\bm{\Phi}_{\mathbb{S}_{3}}(:,q)\|_{2}]^{\bm{\top}}.$
Proof.
Taking $\tau=\gamma=2$ in (4.2), and using (3.4) and (3.3), we obtain
$\mathfrak{K}^{(2),\mathbb{S}}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})=\sup_{\begin{subarray}{c}\left\|\mathrm{v}\mathrm{e}\mathrm{c}\left(\Psi^{\ddagger}\odot\Delta\mathbf{H},\,\bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\right)\right\|_{2}\neq 0\\ {\Delta}A\in\,\mathbb{S}_{1},\,{\Delta}D\in\,\mathbb{S}_{2},\,{\Delta}E\in\,\mathbb{S}_{3}\end{subarray}}\frac{\left\|\Theta_{\xi^{\ddagger}_{\mathrm{L}}}\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\Theta_{\mathrm{v}\mathrm{e}\mathrm{c}(\Psi)}&\bf 0\\ \bf 0&\Theta_{\bm{\chi}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]\right\|_{2}}{\left\|\mathrm{v}\mathrm{e}\mathrm{c}\left(\Psi^{\ddagger}\odot\Delta\mathbf{H},\bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\right)\right\|_{2}}.$
Substituting (4.7) into the above equation yields
$\mathfrak{K}^{(2),\mathbb{S}}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})=\sup_{\begin{subarray}{c}\footnotesize\left\|\left[\begin{matrix}\bm{\Phi}_{\mathbb{S}}&\bf 0\\ \bf 0&I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]\right\|_{2}\neq 0\\ {\Delta}A\in\,\mathbb{S}_{1},\,{\Delta}D\in\,\mathbb{S}_{2},\,{\Delta}E\in\,\mathbb{S}_{3}\end{subarray}}\frac{\left\|\Theta_{\xi^{\ddagger}_{\mathrm{L}}}\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\Theta_{\mathrm{v}\mathrm{e}\mathrm{c}(\Psi)}\bm{\Phi}_{\mathbb{S}}&\bf 0\\ \bf 0&\Theta_{\bm{\chi}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]\right\|_{2}}{\left\|\left[\begin{matrix}\bm{\Phi}_{\mathbb{S}}&\bf 0\\ \bf 0&I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]\right\|_{2}}.$ (4.8)
Utilizing the fact that the matrices $\bm{\Phi}_{\mathbb{S}_{i}}$ for $i=1,2,3,$ are column orthogonal [34], we get $\bm{\Phi}_{\mathbb{S}_{i}}^{\bm{\top}}\bm{\Phi}_{\mathbb{S}_{i}}=\Theta_{\bm{u}_{i}}^{2},$ where $\Theta_{\bm{u}_{i}}$ for $i=1,2,3,$ are the diagonal matrices. Then
$\left\|\left[\begin{matrix}\bm{\Phi}_{\mathbb{S}}&\bf 0\\ \bf 0&I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]\right\|_{2}$ $=\left\|\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]^{\bm{\top}}\left[\begin{matrix}\bm{\Phi}_{\mathbb{S}}^{\bm{\top}}\bm{\Phi}_{\mathbb{S}}&\bf 0\\ \bf 0&I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]\right\|_{2}^{1/2}$
$=\left\|\left[\begin{matrix}\mathfrak{D}_{\mathbb{S}}&\bf 0\\ \bf 0&I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]\right\|_{2}.$ (4.9)
Observe that
$\left[\begin{matrix}\bm{\Phi}_{\mathbb{S}}&\bf 0\\ \bf 0&I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]=\left[\begin{matrix}\bm{\Phi}_{\mathbb{S}}\mathfrak{D}^{-1}_{\mathbb{S}}&\bf 0\\ \bf 0&I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\mathfrak{D}_{\mathbb{S}}&\bf 0\\ \bf 0&I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right].$ (4.10)
Therefore, substituting (4) and (4.10) in (4.8), we obtain
$\mathfrak{K}^{(2),\mathbb{S}}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $=\sup_{\begin{subarray}{c}\footnotesize\left\|\left[\begin{matrix}\mathfrak{D}_{\mathbb{S}}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]\right\|_{2}\neq 0\\ {\Delta}A\in\,\mathbb{S}_{1},\,{\Delta}D\in\,\mathbb{S}_{2},\,{\Delta}E\in\,\mathbb{S}_{3}\end{subarray}}\frac{\left\|\Theta_{\xi^{\ddagger}_{\mathrm{L}}}\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\Theta_{\mathrm{v}\mathrm{e}\mathrm{c}(\Psi)}\bm{\Phi}_{\mathbb{S}}\mathfrak{D}_{\mathbb{S}}^{-1}&\bf 0\\ \bf 0&\Theta_{\bm{\chi}}\end{matrix}\right]\left[\begin{matrix}\mathfrak{D}_{\mathbb{S}}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]\right\|_{2}}{\left\|\left[\begin{matrix}\mathfrak{D}_{\mathbb{S}}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]\right\|_{2}}$
$=\normalsize\left\|\Theta_{\xi^{\ddagger}_{\mathrm{L}}}\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\Theta_{\mathrm{v}\mathrm{e}\mathrm{c}(\Psi)}\bm{\Phi}_{\mathbb{S}}\mathfrak{D}_{\mathbb{S}}^{-1}&\bf 0\\ \bf 0&\Theta_{\bm{\chi}}\end{matrix}\right]\right\|_{2}.$
Hence, the proof is completed. $\blacksquare$
Next, we consider $\tau=\gamma=\infty,$ and derive the structured partial MCN and CCN for the DSPP.
Theorem 5.
Let $A\in\mathbb{S}_{1},$ $D\in\mathbb{S}_{2},$ $E\in\mathbb{S}_{3}$ and $\mathrm{L}\in{\mathbb{R}}^{k\times\bm{l}}.$ Suppose that ${\bm{w}}=[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}$ is the unique solution of DSPP (1.1). When $\tau=\gamma=\infty,$ the structured partial CN of the solution ${\bm{w}}=[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}$ w.r.t. $\mathrm{L}$ is given as follows:
$\mathfrak{K}^{\infty,\,\mathbb{S}}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})=\left\||\Theta_{\xi^{\ddagger}_{\mathrm{L}}}|\left|\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\bm{\Phi}_{\mathbb{S}}&\bf 0\\ \bf 0&I_{\bm{l}}\end{matrix}\right]\right|\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(|\Psi|)\\ |\bm{\chi}|\end{matrix}\right]\right\|_{\infty}.$ (4.11)
Moreover, we consider $\Psi=\mathbf{H},$ $\bm{\chi}={\bf b}.$ Then, if we set $\xi_{\mathrm{L}}=\|\mathrm{L}[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}\|_{\infty}{\bf{1}}_{\bm{l}},$ the structured partial MCN is given by
$\mathfrak{K}^{\infty,\,\mathbb{S}}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})=\frac{\left\|\left|\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\bm{\Phi}_{\mathbb{S}}&\bf 0\\ \bf 0&I_{\bm{l}}\end{matrix}\right]\right|\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(|\mathbf{H}|)\\ |{\bf b}|\end{matrix}\right]\right\|_{\infty}}{\|\mathrm{L}[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}\|_{\infty}}$ (4.12)
and if we set $\xi_{\mathrm{L}}=\mathrm{L}[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}},$ the structured partial CCN is given by
$\mathfrak{K}^{\infty,\,\mathbb{S}}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})=\left\|\frac{\left|\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\bm{\Phi}_{\mathbb{S}}&\bf 0\\ \bf 0&I_{\bm{l}}\end{matrix}\right]\right|\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(|\mathbf{H}|)\\ |{\bf b}|\end{matrix}\right]}{\left|\mathrm{L}[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}\right|}\right\|_{\infty}.$ (4.13)
Proof.
By construction of the matrices $\bm{\Phi}_{\mathbb{S}_{1}},$ $\bm{\Phi}_{\mathbb{S}_{2}}$ and $\bm{\Phi}_{\mathbb{S}_{3}},$ they have almost one nonzero element in each row. Thus, we get
$\left\|\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]\right\|_{\infty}=\left\|\left[\begin{matrix}\bm{\Phi}_{\mathbb{S}}&\bf 0\\ \bf 0&I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]\right\|_{\infty}=\left\|\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]\right\|_{\infty}.$ (4.14)
By considering $\tau=\gamma=\infty$ on (4.2) and using (4.14), (3.4) and (3.3), we obtain
$\mathfrak{K}^{\infty,\mathbb{S}}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $=\sup_{\begin{subarray}{c}\left\|\mathrm{v}\mathrm{e}\mathrm{c}\left(\Psi^{\ddagger}\odot\Delta\mathbf{H},\,\bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\right)\right\|_{\infty}\neq 0\\ {\Delta}A\in\,\mathbb{S}_{1},\,{\Delta}D\in\,\mathbb{S}_{2},\,{\Delta}E\in\,\mathbb{S}_{3}\end{subarray}}\frac{\left\|\Theta_{\xi^{\ddagger}_{\mathrm{L}}}\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\bm{\Phi}_{\mathbb{S}}&\bf 0\\ \bf 0&I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}({\Delta}\mathbf{H})\\ {\Delta}{\bf b}\end{matrix}\right]\right\|_{\infty}}{\left\|\mathrm{v}\mathrm{e}\mathrm{c}\left(\Psi^{\ddagger}\odot\Delta\mathbf{H},\bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\right)\right\|_{\infty}}$
$=\sup_{\footnotesize\begin{subarray}{c}\left\|\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]\right\|_{\infty}\neq 0\\ {\Delta}A\in\,\mathbb{S}_{1},\,{\Delta}D\in\,\mathbb{S}_{2},\,{\Delta}E\in\,\mathbb{S}_{3}\end{subarray}}\frac{\left\|\Theta_{\xi^{\ddagger}_{\mathrm{L}}}\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\bm{\Phi}_{\mathbb{S}}&\bf 0\\ \bf 0&I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\Theta_{\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi)}&\bf 0\\ \bf 0&\Theta_{\bm{\chi}}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]\right\|_{\infty}}{\left\|\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi^{\ddagger}\odot{\Delta}\mathbf{H})\\ \bm{\chi}^{\ddagger}\odot{\Delta}{\bf b}\end{matrix}\right]\right\|_{\infty}}$
$=\left\|\Theta_{\xi^{\ddagger}_{\mathrm{L}}}\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\bm{\Phi}_{\mathbb{S}}&\bf 0\\ \bf 0&I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\Theta_{\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(\Psi)}&\bf 0\\ \bf 0&\Theta_{\bm{\chi}}\end{matrix}\right]\right\|_{\infty}$
$=\left\||\Theta_{\xi^{\ddagger}_{\mathrm{L}}}|\left|\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\mathcal{G}&-I_{\bm{l}}\end{matrix}\right]\left[\begin{matrix}\bm{\Phi}_{\mathbb{S}}&\bf 0\\ \bf 0&I_{\bm{l}}\end{matrix}\right]\right|\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}_{\mathbb{S}}(|\Psi|)\\ |\bm{\chi}|\end{matrix}\right]\right\|_{\infty}.$
The rest of the proof follows by considering $\Psi=\mathbf{H},$ $\bm{\chi}={\bf b},$ $\xi_{\mathrm{L}}=\|\mathrm{L}[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}\|_{\infty}{\bf{1}}_{\bm{l}}$ (or $\xi_{\mathrm{L}}=\mathrm{L}[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}$). $\blacksquare$
5 Deduction of Partial CNs for the EILS problem
The EILS problem is an extension of the famous linear least squares problem, having linear constraints on unknown parameters. It can be expressed as follows:
$\min_{\bm{y}\in{\mathbb{R}}^{m}}(b-M\bm{y})^{\bm{\top}}\mathbb{J}\leavevmode\nobreak\ (b-M\bm{y})\quad\text{subjected to}\quad C\bm{y}=d,$ (5.1)
where $M\in{\mathbb{R}}^{n\times m}(n\geq m),$ $C\in{\mathbb{R}}^{p\times m},$ $b\in{\mathbb{R}}^{n}$ $d\in{\mathbb{R}}^{p}$ and the signature matrix $\mathbb{J}$ given by
$\mathbb{J}=\left[\begin{matrix}I_{n_{1}}&\bf 0\\ \bf 0&-I_{n_{2}}\end{matrix}\right],\leavevmode\nobreak\ n_{1}+n_{2}=n.$ (5.2)
When $\mathrm{rank}(C)=p$ and $\bm{y}^{\bm{\top}}(M^{\bm{\top}}\mathbb{J}M)\bm{y}>0$ for all nonzero $\bm{y}\in\text{null}(C),$ the EILS problem (5.1) has a unique solution. The solution of the EILS (5.1) problem also satisfies the following the augmented system [24]:
$\widehat{\mathfrak{B}}\left[\begin{matrix}\lambda\\ \bm{x}\\ \bm{y}\end{matrix}\right]:=\left[\begin{matrix}\bf 0&\bf 0&C\\ \bf 0&\mathbb{J}&M\\ C^{\bm{\top}}&M^{\bm{\top}}&\bf 0\end{matrix}\right]\left[\begin{matrix}\lambda\\ \bm{x}\\ \bm{y}\end{matrix}\right]=\left[\begin{matrix}d\\ b\\ \bf 0\end{matrix}\right],$ (5.3)
where $\bm{x}=\mathbb{J}{\bm{r}},$ ${\bm{r}}=b-M\bm{y}$ and $\lambda=(CC^{\bm{\top}})^{-1}CM^{\bm{\top}}\mathbb{J}{\bm{r}}$ is the vector of Lagrange multipliers [35]. Note that the system in (5.3) can be equivalently transformed into
$\left[\begin{matrix}\mathbb{J}&M&\bf 0\\ M^{\bm{\top}}&\bf 0&C^{\bm{\top}}\\ \bf 0&C&\bf 0\end{matrix}\right]\left[\begin{matrix}\bm{x}\\ \bm{y}\\ \lambda\end{matrix}\right]=\left[\begin{matrix}b\\ \bf 0\\ d\end{matrix}\right]=:\widehat{\bf d}.$ (5.4)
Observe that, the above system is in the form of DSPP (1.1) with $A=\mathbb{J},$ $B=M^{\bm{\top}},$ ${\bf b}=[b^{\bm{\top}},{\bf 0},d^{\bm{\top}}]^{\bm{\top}}$ and $\bm{z}=\lambda.$ Therefore, the task of assessing the conditioning of the EILS problem (5.1) can be achieved by determining the CNs for the solution $\bm{y}$ of DSPP (5.4).
Generally, the signature matrix $\mathbb{J}$ has no perturbation and as $D={\bf 0},$ $E={\bf 0}$ and ${\bf b}_{2}={\bf 0}$, we consider ${\Delta}A=\bf 0$, ${\Delta}D=\bf 0,$ ${\Delta}E=0$ and ${\Delta}{\bf b}_{2}={\bf 0}$ in (2.12). Then, the perturbation expression in (2.13) reduces to
$\left[\begin{matrix}{\Delta}\bm{x}\\ {\Delta}\bm{z}\\ {\Delta}\bm{y}\end{matrix}\right]=-\mathfrak{B}^{-1}\left[\begin{matrix}\widehat{\mathcal{G}}&-I_{n+p}\end{matrix}\right]\left[\begin{matrix}\mathrm{v}\mathrm{e}\mathrm{c}({\Delta}B^{{\bm{\top}}})\\ \mathrm{v}\mathrm{e}\mathrm{c}({\Delta}C)\\ {\Delta}{\bf b}_{1}\\ {\Delta}{\bf b}_{3}\end{matrix}\right]+\mathcal{O}(\bm{\epsilon}^{2}),$ (5.5)
where
$\widehat{\mathcal{G}}=\left[\begin{matrix}\bm{y}^{\bm{\top}}\otimes I_{n}&\bf 0\\ I_{m}\otimes\bm{x}^{\bm{\top}}&I_{m}\otimes\bm{z}^{\bm{\top}}\\ \bf 0&\bm{y}^{\bm{\top}}\otimes I_{p}\end{matrix}\right]\in{\mathbb{R}}^{\bm{l}\times\hat{\bm{s}}},$ (5.6)
$\hat{\bm{s}}=m(n+p).$
Let $\widehat{\mathbf{H}}=(B^{{\bm{\top}}},C),$ ${\Delta}\widehat{\mathbf{H}}=({\Delta}B^{{\bm{\top}}},{\Delta}C),$ $\widehat{\bf b}=[{\bf b}_{1}^{\bm{\top}},{\bf b}_{3}^{\bm{\top}}]^{\bm{\top}}$ and ${\Delta}\widehat{\bf b}=[{\Delta}{\bf b}_{1}^{\bm{\top}},{\Delta}{\bf b}_{3}^{\bm{\top}}]^{\bm{\top}}$. We define the following mapping:
$\widetilde{\bm{\varphi}}:\,$ ${\mathbb{R}}^{m\times n}\times{\mathbb{R}}^{p\times m}\times{\mathbb{R}}^{n+p}\rightarrow{\mathbb{R}}^{k}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \widetilde{\bm{\varphi}}(\widehat{\mathbf{H}},\widehat{\mathbf{b}})=\mathrm{L}\left[\begin{matrix}\bm{x}\\ \bm{y}\\ \bm{z}\end{matrix}\right]=\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}{\bf b}_{1}\\ \bf 0\\ {\bf b}_{3}\end{matrix}\right],$ (5.7)
where $\mathrm{L}\in{\mathbb{R}}^{k\times\bm{l}}.$ Using a similar method to the Theorem 1, we have the following result.
Theorem 6.
Assume that $[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}$ is the unique solution of the DSPP (1.1) with $D=\bf 0,$ $E=\bf 0$, ${\bf b}_{2}={\bf 0}$ and $\mathrm{L}\in{\mathbb{R}}^{k\times\bm{l}}.$ Then the partial unified CN of $[\bm{x}^{\bm{\top}},\bm{y}^{\bm{\top}},\bm{z}^{\bm{\top}}]^{\bm{\top}}$ w.r.t. $\mathrm{L}$ is given by
$\mathfrak{K}_{\widetilde{\bm{\varphi}}}(\widehat{\mathbf{H}},{\bf b};\mathrm{L})$ $:=\lim_{\bm{\epsilon}\rightarrow 0}\sup_{0<\left\|\mathrm{v}\mathrm{e}\mathrm{c}\left(\widehat{\Psi}^{\ddagger}\odot\Delta\widehat{\mathbf{H}},\,\widehat{\bm{\chi}}^{\ddagger}\odot{\Delta}{\bf b}\right)\right\|_{\tau}\leq\bm{\epsilon}}\frac{\left\|\xi^{\ddagger}_{\mathrm{L}}\odot\left(\widetilde{\bm{\varphi}}(\widehat{\mathbf{H}}+{\Delta}\widehat{\mathbf{H}},\widehat{\mathbf{b}}+{\Delta}\widehat{\mathbf{b}})-\widetilde{\bm{\varphi}}(\widehat{\mathbf{H}},\widehat{{\bf b}})\right)\right\|_{\gamma}}{\left\|\mathrm{v}\mathrm{e}\mathrm{c}\left(\widehat{\Psi}^{\ddagger}\odot\Delta\widehat{\mathbf{H}},\widehat{\bm{\chi}}^{\ddagger}\odot{\Delta}\widehat{{\bf b}}\right)\right\|_{\tau}},$
$\,=\left\|\Theta_{\xi^{\ddagger}_{\mathrm{L}}}\mathrm{L}\mathfrak{B}^{-1}\left[\begin{matrix}\widehat{\mathcal{G}}&-I_{n+p}\end{matrix}\right]\left[\begin{matrix}\Theta_{\mathrm{v}\mathrm{e}\mathrm{c}(\widehat{\Psi})}&\bf 0\\ \bf 0&\Theta_{\widehat{\bm{\chi}}}\end{matrix}\right]\right\|_{\tau,\gamma},$
where $\widehat{\Psi}=(\Psi_{B^{{\bm{\top}}}},\Psi_{C}),$ $\Psi_{B^{{\bm{\top}}}}\in{\mathbb{R}}^{n\times m},\Psi_{C}\in{\mathbb{R}}^{p\times m}$ and $\widehat{\bm{\chi}}\in{\mathbb{R}}^{n+p}.$
Remark 6.
[…]
$\Sigma=\left[\begin{matrix}\bf 0&I_{n}&\bf 0\\ \bf 0&\bf 0&I_{m}\\ I_{p}&\bf 0&\bf 0\end{matrix}\right],$ (5.8)
using the formula for the inverse of $\widehat{\mathfrak{B}}$ given in [36], we obtain
$\mathfrak{K}_{\widetilde{\bm{\varphi}}}(\widehat{\mathbf{H}},{\bf b};\mathrm{L})=\left\|\Theta_{\xi^{\ddagger}_{\mathrm{L}}}\mathrm{L}_{1}\left[\begin{matrix}\Xi&\Lambda&-(\mathcal{Q}\mathcal{M}\mathcal{Q})^{\dagger}MJ&\mathcal{B}_{M}\end{matrix}\right]\left[\begin{matrix}\Theta_{\mathrm{v}\mathrm{e}\mathrm{c}(\widehat{\Psi})}&\bf 0\\ \bf 0&\Theta_{\widehat{\bm{\chi}}}\end{matrix}\right]\right\|_{\tau,\gamma},$ (5.9)
where $\Xi=\bm{y}^{\bm{\top}}\otimes(\mathcal{Q}\mathcal{M}\mathcal{Q})^{\dagger}M\mathbb{J}-(\mathcal{Q}\mathcal{M}\mathcal{Q})^{\dagger}\otimes\bm{x}^{\bm{\top}},$ $\Lambda=\bm{y}^{\bm{\top}}\otimes\mathcal{B}_{M}-(\mathcal{Q}\mathcal{M}\mathcal{Q})^{\dagger}\otimes\bm{z}^{\bm{\top}},$ $\mathcal{M}=M\mathbb{J}M^{\bm{\top}},$ $\mathcal{Q}=I_{m}-C^{\dagger}C$ and $\mathcal{B}_{M}=(I_{m}-\mathcal{Q}\mathcal{M}\mathcal{Q})^{\dagger}.$ Note that the partial CN expression is the same as derived in [24].
6 Numerical experiments
In this part, we present some numerical examples to verify the reliability of the derived the partial NCN, MCN and CCN and their upper bounds for the DSPP. Additionally, we demonstrate their effectiveness in providing tight upper bounds for the relative forward error in solving the DSPP.
We construct the entrywise perturbation as follows:
${\Delta}A=10^{-s}\cdot\mathtt{randn}(n,n)\odot A,\leavevmode\nobreak\ {\Delta}B=10^{-s}\cdot\mathtt{randn}(m,n)\odot B,\leavevmode\nobreak\ {\Delta}C=10^{-s}\cdot\mathtt{randn}(p,m)\odot C,$
${\Delta}D=10^{-s}\cdot\mathtt{randn}(m,m)\odot D,\leavevmode\nobreak\ {\Delta}E=10^{-s}\cdot\mathtt{randn}(p,p)\odot E,\leavevmode\nobreak\ {\Delta}{\bf b}=10^{-s}\cdot\mathtt{randn}(\bm{l},1)\odot{\bf b},$
where $\mathtt{randn}(m,n)$ denotes $m\times n$ the random matrices generated by MATLAB command $\mathtt{randn}$. Let $\bm{w}=[\bm{x}^{{\bm{\top}}},\bm{y}^{{\bm{\top}}},\bm{z}^{\bm{\top}}]^{{\bm{\top}}}$ and $\widetilde{\bm{w}}=[\widetilde{\bm{x}}^{{\bm{\top}}},\widetilde{\bm{y}}^{{\bm{\top}}},\widetilde{\bm{z}}^{{\bm{\top}}}]^{{\bm{\top}}}$ be the unique solutions of the original DSPP and the perturbed DSPP, respectively. To estimate an upper bound for the forward error in the solution, the normwise, mixed and componentwise relative forward errors are defined as follows:
$\bm{r}_{k}=\frac{\|\mathrm{L}\widetilde{\bm{w}}-\mathrm{L}\bm{w}\|_{2}}{\|\mathrm{L}\bm{w}\|_{2}},\quad\bm{r}_{m}=\frac{\|\mathrm{L}\widetilde{\bm{w}}-\mathrm{L}\bm{w}\|_{\infty}}{\|\mathrm{L}\bm{w}\|_{\infty}}\quad\text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \bm{r}_{c}=\left\|\frac{\mathrm{L}\widetilde{\bm{w}}-\mathrm{L}\bm{w}}{\mathrm{L}\bm{w}}\right\|_{\infty},$
respectively. Define the following quantities:
$\bm{\epsilon}_{1}=\frac{\left\|\left[\begin{matrix}{\Delta}\mathfrak{B}&{\Delta}\mathbf{b}\end{matrix}\right]\right\|_{F}}{\left\|\left[\begin{matrix}\mathfrak{B}&\mathbf{b}\end{matrix}\right]\right\|_{F}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \bm{\epsilon}_{2}=\min\left\{\bm{\epsilon}:\left|{\Delta}\mathfrak{B}\right|\leq\bm{\epsilon}\left|\mathfrak{B}\right|,\leavevmode\nobreak\ \left|{\Delta}\mathbf{b}\right|\leq\bm{\epsilon}\left|\mathbf{b}\right|\right\}.$ (6.1)
Thus from (2.11), $\bm{\epsilon}_{1}\mathfrak{K}^{(2)}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}),$ $\bm{\epsilon}_{2}\mathfrak{K}^{\infty}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ and $\bm{\epsilon}_{2}\mathfrak{K}^{\infty}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ can be employed to estimate the relative forward errors $\bm{r}_{k},$ $\bm{r}_{m}$ and $\bm{r}_{c},$ respectively. We select the matrix $\mathrm{L}$ as
$\mathrm{L}_{0}=I_{\bm{l}},\leavevmode\nobreak\ \mathrm{L}_{n}=\left[\begin{matrix}I_{n}&\mathbf{0}_{n\times(m+p)}\end{matrix}\right],\leavevmode\nobreak\ \mathrm{L}_{m}=\left[\begin{matrix}{\bf 0}_{m\times n}&I_{m}&\mathbf{0}_{m\times p}\end{matrix}\right]\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \mathrm{L}_{p}=\left[\begin{matrix}{\bf 0}_{p\times(n+m)}&I_{p}\end{matrix}\right]$
to obtain the CNs for $\bm{w}$, $\bm{x}$, $\bm{y}$ and $\bm{z}$, respectively.
Example 6.1.
We consider the DSPP (1.1) taken from [8] with
$A=\left[\begin{matrix}I_{q}\otimes J+J\otimes I_{q}&\bf 0\\ \bf 0&I_{q}\otimes J+J\otimes I_{q}\end{matrix}\right]\in{\mathbb{R}}^{2q^{2}\times 2q^{2}},\leavevmode\nobreak\ \leavevmode\nobreak$
$B=\left[\begin{matrix}I_{q}\otimes Z&Z\otimes I_{q}\end{matrix}\right]\in{\mathbb{R}}^{q^{2}\times 2q^{2}},\leavevmode\nobreak\ C=Y\otimes Z\in{\mathbb{R}}^{q^{2}\times q^{2}},$
where $J=\frac{1}{(q+1)^{2}}\,\mathtt{tridiag}(-1,2,-1)\in{\mathbb{R}}^{q\times q},$ $Z=\frac{1}{q+1}\,\mathtt{tridiag}(0,1,-1)\in{\mathbb{R}}^{q\times q}$ and $Y=\mathrm{diag}(1,q+1,\ldots,q^{2}-q+1)\in{\mathbb{R}}^{q\times q}.$ Further, we take $D=I_{m}$ and $E=I_{p}.$ The notation $\mathtt{tridiag}(d_{1},d_{2},d_{3})\in{\mathbb{R}}^{q\times q}$ denotes the tridiagonal matrix with subdiagonal entries $d_{1}$, diagonal entries $d_{2}$ and superdiagonal entries $d_{3}$. Here, the dimension of the coefficient matrix $\mathfrak{B}$ of the DSPP is ${\bm{l}}=4q^{2}.$ We take ${\bf b}=\tt{randn}(n,1)\in{\mathbb{R}}^{\bm{l}}$. Further, we consider $\Psi=\|\mathfrak{B}\|_{F}$ and ${\bm{\chi}}=\|{\bf b}\|_{2}.$ We use Theorems 2, 3 to compute partial CNs and Corollaries 3.1 and 3.2 for their upper bounds. The numerical results for different choices for $\mathrm{L}$ and $q=4:2:16$ are presented in Tables 1-4.
Table 1: Comparison of the NCN, MCN, and CCN, and their upper bounds, with the corresponding relative errors for $\mathrm{L}=\mathrm{L}_{0}$ for Example 6.1.
$q$ $\bm{r}_{k}$ $\mathfrak{K}_{2}$ $\mathfrak{K}_{2}^{U}$ $\bm{r}_{m}$ $\mathfrak{K}_{m}$ $\mathfrak{K}_{m}^{U}$ $\bm{r}_{c}$ $\mathfrak{K}_{c}$ $\mathfrak{K}_{c}^{U}$
4 $081.1882E-$ 08 $065.1234E-$ 06 $066.3888E-$ 06 $081.9907E-$ 08 $073.5552E-$ 07 $073.8637E-$ 07 $072.1532E-$ 07 $066.9184E-$ 06 $068.6300E-$ 06
$6$ $082.0426E-$ 08 $068.2886E-$ 06 $051.6984E-$ 05 $082.3603E-$ 08 $073.9780E-$ 07 $074.0885E-$ 07 $073.7248E-$ 07 $051.2135E-$ 05 $051.2490E-$ 05
$8$ $083.1951E-$ 08 $051.9154E-$ 05 $053.6586E-$ 05 $084.1045E-$ 08 $074.9025E-$ 07 $074.9808E-$ 07 $068.3715E-$ 06 $041.3746E-$ 04 $041.3964E-$ 04
$10$ $082.9187E-$ 08 $055.5170E-$ 05 $041.1519E-$ 04 $084.2771E-$ 08 $061.2319E-$ 06 $061.2706E-$ 06 $071.8721E-$ 07 $051.9213E-$ 05 $052.1515E-$ 05
$12$ $082.3289E-$ 08 $059.4616E-$ 05 $041.7034E-$ 04 $082.9289E-$ 08 $061.0722E-$ 06 $061.0798E-$ 06 $074.5669E-$ 07 $053.5850E-$ 05 $053.8136E-$ 05
$14$ $083.9431E-$ 08 $041.6571E-$ 04 $043.3889E-$ 04 $083.9739E-$ 08 $061.2469E-$ 06 $061.2747E-$ 06 $076.0883E-$ 07 $053.1225E-$ 05 $053.5558E-$ 05
$16$ $083.9038E-$ 08 $042.7496E-$ 04 $045.6909E-$ 04 $083.9932E-$ 08 $061.8001E-$ 06 $061.8356E-$ 06 $064.2789E-$ 06 $041.7386E-$ 04 $041.8883E-$ 04
Table 2: Comparison of the NCN, MCN, and CCN, and their upper bounds, with the corresponding relative errors for $\mathrm{L}=\mathrm{L}_{n}$ for Example 6.1.
$q$ $\bm{r}_{k}$ $\mathfrak{K}_{2}$ $\mathfrak{K}_{2}^{U}$ $\bm{r}_{m}$ $\mathfrak{K}_{m}$ $\mathfrak{K}_{m}^{U}$ $\bm{r}_{c}$ $\mathfrak{K}_{c}$ $\mathfrak{K}_{c}^{U}$
$4$ $082.3210E-$ 08 $061.7899E-$ 06 $062.2325E-$ 06 $084.4418E-$ 08 $073.9328E-$ 07 $073.9952E-$ 07 $086.1112E-$ 08 $077.7010E-$ 07 $077.9196E-$ 07
$6$ $082.2133E-$ 08 $062.6980E-$ 06 $067.6732E-$ 06 $084.2425E-$ 08 $078.8875E-$ 07 $079.5134E-$ 07 $085.3918E-$ 08 $065.8823E-$ 06 $067.3855E-$ 06
$8$ $083.8470E-$ 08 $064.8135E-$ 06 $051.5701E-$ 05 $088.7676E-$ 08 $061.4710E-$ 06 $061.5322E-$ 06 $075.7409E-$ 07 $055.8895E-$ 05 $057.0437E-$ 05
$10$ $082.9894E-$ 08 $069.6923E-$ 06 $053.1204E-$ 05 $085.9374E-$ 08 $062.0857E-$ 06 $062.1065E-$ 06 $072.7350E-$ 07 $059.5571E-$ 05 $041.0598E-$ 04
$12$ $083.3113E-$ 08 $053.6464E-$ 05 $054.9737E-$ 05 $085.6822E-$ 08 $063.0914E-$ 06 $063.0928E-$ 06 $074.5080E-$ 07 $069.3747E-$ 06 $069.4743E-$ 06
$14$ $088.0906E-$ 08 $052.4824E-$ 05 $057.6132E-$ 05 $071.5910E-$ 07 $063.3988E-$ 06 $063.4220E-$ 06 $078.4585E-$ 07 $052.5939E-$ 05 $052.9725E-$ 05
$16$ $084.4663E-$ 08 $053.5262E-$ 05 $058.8133E-$ 05 $088.0314E-$ 08 $064.7189E-$ 06 $064.7301E-$ 06 $074.5566E-$ 07 $053.7542E-$ 05 $053.9779E-$ 05
Table 3: Comparison of the NCN, MCN, and CCN, and their upper bounds, with the corresponding relative errors for $\mathrm{L}=\mathrm{L}_{m}$ for Example 6.1.
$q$ $\bm{r}_{k}$ $\mathfrak{K}_{2}$ $\mathfrak{K}_{2}^{U}$ $\bm{r}_{m}$ $\mathfrak{K}_{m}$ $\mathfrak{K}_{m}^{U}$ $\bm{r}_{c}$ $\mathfrak{K}_{c}$ $\mathfrak{K}_{c}^{U}$
$4$ $081.4640E-$ 08 $064.1828E-$ 06 $068.1301E-$ 06 $081.7030E-$ 08 $073.2527E-$ 07 $073.6699E-$ 07 $086.3575E-$ 08 $061.6374E-$ 06 $061.7734E-$ 06
$6$ $083.8057E-$ 08 $053.4032E-$ 05 $055.4748E-$ 05 $083.9292E-$ 08 $078.5355E-$ 07 $061.0120E-$ 06 $086.8320E-$ 08 $062.2698E-$ 06 $062.3143E-$ 06
$8$ $086.2425E-$ 08 $041.1653E-$ 04 $041.7568E-$ 04 $085.3397E-$ 08 $061.9931E-$ 06 $062.1774E-$ 06 $073.0610E-$ 07 $051.3012E-$ 05 $051.3193E-$ 05
$10$ $084.7836E-$ 08 $041.5526E-$ 04 $043.3655E-$ 04 $085.1330E-$ 08 $061.8322E-$ 06 $062.2332E-$ 06 $075.0917E-$ 07 $052.8394E-$ 05 $052.8957E-$ 05
$12$ $083.7645E-$ 08 $059.3405E-$ 05 $044.7395E-$ 04 $083.6702E-$ 08 $061.1129E-$ 06 $061.4824E-$ 06 $087.7646E-$ 08 $067.9413E-$ 06 $068.8422E-$ 06
$14$ $087.7970E-$ 08 $041.6058E-$ 04 $049.2172E-$ 04 $085.7666E-$ 08 $061.7296E-$ 06 $062.1704E-$ 06 $062.9851E-$ 06 $057.6483E-$ 05 $057.9562E-$ 05
$16$ $084.7259E-$ 08 $044.3435E-$ 04 $031.5265E-$ 03 $086.6721E-$ 08 $062.2484E-$ 06 $062.8853E-$ 06 $072.9895E-$ 07 $055.8651E-$ 05 $056.0528E-$ 05
Table 4: Comparison of the NCN, MCN, and CCN, and their upper bounds, with the corresponding relative errors for $\mathrm{L}=\mathrm{L}_{p}$ for Example 6.1.
$q$ $\bm{r}_{k}$ $\mathfrak{K}_{2}$ $\mathfrak{K}_{2}^{U}$ $\bm{r}_{m}$ $\mathfrak{K}_{m}$ $\mathfrak{K}_{m}^{U}$ $\bm{r}_{c}$ $\mathfrak{K}_{c}$ $\mathfrak{K}_{c}^{U}$
$4$ $097.9587E-$ 09 $063.2626E-$ 06 $065.0845E-$ 06 $099.2868E-$ 09 $072.7450E-$ 07 $072.8678E-$ 07 $086.4428E-$ 08 $075.9782E-$ 07 $076.2833E-$ 07
$6$ $081.3913E-$ 08 $069.6919E-$ 06 $051.1609E-$ 05 $081.8364E-$ 08 $074.4130E-$ 07 $074.4401E-$ 07 $084.1485E-$ 08 $061.1621E-$ 06 $061.1751E-$ 06
$8$ $082.1543E-$ 08 $052.5173E-$ 05 $054.6780E-$ 05 $082.7719E-$ 08 $077.9552E-$ 07 $078.4231E-$ 07 $084.9306E-$ 08 $064.2242E-$ 06 $064.4848E-$ 06
$10$ $083.2132E-$ 08 $055.7696E-$ 05 $041.1250E-$ 04 $084.3437E-$ 08 $061.0618E-$ 06 $061.1161E-$ 06 $088.3724E-$ 08 $064.1932E-$ 06 $064.3258E-$ 06
$12$ $081.6735E-$ 08 $058.7652E-$ 05 $041.4248E-$ 04 $083.1936E-$ 08 $061.0011E-$ 06 $061.0138E-$ 06 $086.2623E-$ 08 $062.5824E-$ 06 $062.5860E-$ 06
$14$ $082.0065E-$ 08 $041.5497E-$ 04 $041.9181E-$ 04 $082.2200E-$ 08 $061.0421E-$ 06 $061.0457E-$ 06 $087.7378E-$ 08 $064.1703E-$ 06 $064.1741E-$ 06
$16$ $081.0902E-$ 08 $042.4487E-$ 04 $043.5342E-$ 04 $081.5817E-$ 08 $061.6369E-$ 06 $061.6456E-$ 06 $071.0229E-$ 07 $063.4527E-$ 06 $063.4614E-$ 06
For simplicity, we denote:
$\mathfrak{K}_{2}=\bm{\epsilon}_{1}\mathfrak{K}^{(2)}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}_{m}),\leavevmode\nobreak\ \mathfrak{K}_{2}^{U}=\bm{\epsilon}_{1}\mathfrak{K}^{(2),u}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}_{m}),\leavevmode\nobreak\ \mathfrak{K}_{m}=\bm{\epsilon}_{2}\mathfrak{K}^{\infty}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}_{m}),$
$\mathfrak{K}_{m}^{U}=\bm{\epsilon}_{2}\mathfrak{K}^{\infty,u}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}_{m}),\leavevmode\nobreak\ \mathfrak{K}_{c}=\bm{\epsilon}_{2}\mathfrak{K}^{\infty}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}_{m}),\leavevmode\nobreak\ \mathfrak{K}_{c}^{U}=\bm{\epsilon}_{2}\mathfrak{K}^{\infty,u}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}_{m}).$
The results in Tables 1-4 show that the upper bounds of the NCN, MCN and CCN provide very sharp estimates of the exact NCN, MCN and CCN, respectively. Additionally, it is observed that both the MCN and CCN, along with their upper bounds, are at most two orders of magnitude larger than the actual relative forward errors, offering more accurate estimates compared to the NCN and its upper bounds. These numerical results highlight the effectiveness of the proposed CNs and their corresponding upper bounds.
Example 6.2.
We consider the DSPP (1.1) with the block matrices given by
$A=[(2ZZ^{T}+\Sigma_{1})\oplus\Sigma_{2}\oplus\Sigma_{3}]\in{\mathbb{R}}^{n\times n},\leavevmode\nobreak\ B=\left[\begin{matrix}N&-I_{2\tilde{q}}&I_{2\tilde{q}}\end{matrix}\right]\in{\mathbb{R}}^{m\times n},\leavevmode\nobreak$
$D={\tt{toeplitz}}(\bm{d})\in{\mathbb{R}}^{m\times m},\leavevmode\nobreak\ C=M,\leavevmode\nobreak\ E={\tt{toeplitz}}(\bm{e})\in{\mathbb{R}}^{p\times p},$
$\leavevmode\nobreak\ \text{where}\leavevmode\nobreak\ Z=[z_{ij}]\in{\mathbb{R}}^{\hat{q}\times\hat{q}}\leavevmode\nobreak\ \text{with}\leavevmode\nobreak\ z_{ij}=e^{-2((i/3)^{2}+(j/3)^{2})},\leavevmode\nobreak\ \Sigma_{1}=I_{\hat{q}},$
and $\Sigma_{k}=\mathrm{diag}(d_{j}^{(k)})\in{\mathbb{R}}^{2\tilde{q}\times 2\tilde{q}},\leavevmode\nobreak\ k=2,3,$ are diagonal matrices with
$d_{j}^{(2)}=\left\{\begin{array}[]{cc}1,&\text{for}\leavevmode\nobreak\ 1\leq j\leq\tilde{q},\\ 10^{-5}(j-\tilde{q})^{2},&\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \tilde{q}+1\leq j\leq 2\tilde{q},\end{array}\right.$
$d^{(3)}_{j}=10^{-5}(j+\tilde{q})^{2}$ for $1\leq j\leq 2\tilde{q},$ where $\tilde{q}=q^{2}$ and $\hat{q}=q(q+1).$ Further, $N=\left[\begin{matrix}\widehat{N}\otimes I_{q}\\ I_{q}\otimes\widehat{N}\end{matrix}\right]\in{\mathbb{R}}^{2\tilde{q}\times\hat{q}},$ $\widehat{N}={\tt tridiag}(0,2,-1)\in{\mathbb{R}}^{q\times(q+1)},$ $M=\left[\begin{matrix}\widehat{M}\otimes I_{q}&I_{q}\otimes\widehat{M}\end{matrix}\right]\in{\mathbb{R}}^{\hat{q}\times 2\tilde{q}},$
$\widehat{M}=\left[\begin{matrix}q+1&-\frac{q-1}{q}&\frac{q-2}{q}&\ldots&\frac{(-1)^{q-1}}{q}\\ -\frac{q-1}{q}&2q+1&\ddots&\ddots&\vdots\\ \frac{q-2}{q}&\ddots&\ddots&\ddots&\frac{q-2}{q}\\ \vdots&\ddots&\ddots&\ddots&-\frac{q-1}{q}\\ \frac{(-1)^{q-1}}{q}&\vdots&\frac{q-2}{q}&-\frac{q-1}{q}&q^{2}+1\\ 0&0&\ldots&0&1\end{matrix}\right]\in{\mathbb{R}}^{(q+1)\times q},$
$\bm{d}=\mathtt{randn}(m,1)\in{\mathbb{R}}^{m},\bm{e}=\mathtt{randn}(p,1)\in{\mathbb{R}}^{p}.$ Here, the notation $[A\oplus B]$ means the block diagonal matrix with diagonal blocks $A$ and $B,$ and $\mathtt{toeplitz}(\bm{d})$ denotes the Toeplitz matrix generated by the vector $\bm{d}.$ Moreover, we get ${\bm{l}}=8q^{2}+2q.$ The vector ${\bf d}\in{\mathbb{R}}^{\bm{l}}$ is chosen as in Example 6.1.
Table 5: Comparison of the partial NCN, MCN, and CCN with the corresponding structured partial NCN, MCN and CCN for Example 6.2.
$\mathrm{L}$ $q$ $\mathfrak{K}^{(2)}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $\mathfrak{K}^{(2),\mathbb{S}}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $\mathfrak{K}^{\infty}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $\mathfrak{K}^{\infty,\mathbb{S}}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $\mathfrak{K}^{\infty}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $\mathfrak{K}^{\infty,\mathbb{S}}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$
$\mathrm{L}_{0}$ $2$ $039.0961E+$ 03 $038.9842E+$ 03 $021.1174E+$ 02 $019.2003E+$ 01 $031.0211E+$ 03 $027.8013E+$ 02
$3$ $041.0996E+$ 04 $038.6963E+$ 03 $021.1735E+$ 02 $019.1218E+$ 01 $041.2372E+$ 04 $039.5905E+$ 03
$4$ $042.6066E+$ 04 $041.9983E+$ 04 $023.5456E+$ 02 $022.6052E+$ 02 $052.5284E+$ 05 $046.7124E+$ 04
$\mathrm{L}_{n}$ $2$ $039.1011E+$ 03 $038.9891E+$ 03 $021.1174E+$ 02 $019.2003E+$ 01 $022.9170E+$ 02 $022.1175E+$ 02
$3$ $041.0972E+$ 04 $038.6781E+$ 03 $021.1735E+$ 02 $019.1218E+$ 01 $041.2372E+$ 04 $039.5905E+$ 03
$4$ $042.6371E+$ 04 $042.0138E+$ 04 $023.5456E+$ 02 $022.6052E+$ 02 $042.7573E+$ 04 $041.8454E+$ 04
$\mathrm{L}_{m}$ $2$ $038.4606E+$ 03 $038.3300E+$ 03 $021.5279E+$ 02 $021.1486E+$ 02 $022.5276E+$ 02 $021.5433E+$ 02
$3$ $041.8511E+$ 04 $041.4614E+$ 04 $022.2517E+$ 02 $021.6131E+$ 02 $029.9621E+$ 02 $025.3614E+$ 02
$4$ $043.2837E+$ 04 $042.7473E+$ 04 $028.9167E+$ 02 5.8061E+02 $052.5284E+$ 05 $046.7124E+$ 04
$\mathrm{L}_{p}$ $2$ $041.1952E+$ 04 $041.1786E+$ 04 $022.1568E+$ 02 $021.6042E+$ 02 $031.0211E+$ 03 $027.8013E+$ 02
$3$ $039.4332E+$ 03 $037.4505E+$ 03 $021.2141E+$ 02 $018.3080E+$ 01 $041.1802E+$ 04 $038.2273E+$ 03
$4$ $041.8803E+$ 04 $041.5754E+$ 04 $027.0206E+$ 02 $024.8744E+$ 02 $037.4731E+$ 03 $034.8946E+$ 03
The structured partial NCN is calculated using Theorem 4, while the structured partial MCN and CCN are computed from Theorem 5. Numerical results for various choices of $\mathrm{L}$ and $q=2,3,4$ are summarized in Table 5. We observed for all choices of $\mathrm{L},$ $\mathfrak{K}^{\infty,\mathbb{S}}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ and $\mathfrak{K}^{\infty,\mathbb{S}}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ are almost one order smaller than the $\mathfrak{K}^{\infty}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ and $\mathfrak{K}^{\infty}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$, respectively.
7 Conclusion
This paper introduces a unified framework for investigating the partial CN for the solution of the DSPP. We derive compact formulas for the partial CNs, and by considering specific norms, we obtain the partial NCN, MCN and CCN for the solution of the DSPP. Additionally, we provide sharp upper bounds for the partial CNs that are free from costly Kronecker products. Moreover, we compute structured partial NCN, MCN, and CCN by introducing perturbations that maintain the structure of the block matrices of the coefficient matrix. Using our theoretical findings and by leveraging the relationship between the EILS problems and the DSPP, we recover several previously established results for the EILS problems. Experimental results demonstrate that the derived upper bounds for the partial CNs provide tight estimates of the actual partial CNs. Furthermore, the proposed partial CNs and their upper bounds provide sharp error estimation for the solution, highlighting their effectiveness and reliability.
[…]
Acknowledgements During this work, Pinki Khatun was financially supported by the Council of Scientific &\&& Industrial Research (CSIR) in New Delhi, India, in the form of a fellowship (File no. $09/1022(0098)/2020$-EMR-I).
Declarations
Funding
Not applicable
Competing interests
The authors have no competing interests
Ethics approval
Not applicable
Data availability
Not applicable
Materials availability
Not applicable
Author contribution
Both authors contributed equally to the conceptualization, writing, and review of this work.